### v1

# An End-to-End Room Geometry Constrained Depth Estimation Framework for Indoor Panorama Images

Abstract
Predicting spherical pixel depth from monocular $360^{\circ}$ indoor panoramas is critical for many vision applications. However, existing methods focus on pixel-level accuracy, causing oversmoothed room corners and noise sensitivity. In this paper, we propose a depth estimation framework based on room geometry constraints, which extracts room geometry information through layout prediction and integrates those information into the depth estimation process through background segmentation mechanism. At the model level, our framework comprises a shared feature encoder followed by task-specific decoders for layout estimation, depth estimation, and background segmentation. The shared encoder extracts multi-scale features, which are subsequently processed by individual decoders to generate initial predictions: a depth map, a room layout map, and a background segmentation map. Furthermore, our framework incorporates two strategies: a room geometry-based background depth resolving strategy and a background-segmentation-guided fusion mechanism. The proposed room-geometry-based background depth resolving strategy leverages the room layout and the depth decoder’s output to generate the corresponding background depth map. Then, a background-segmentation-guided fusion strategy derives fusion weights for the background and coarse depth maps from the segmentation decoder’s predictions. Extensive experimental results on the Stanford2D3D, Matterport3D and Structured3D datasets show that our proposed methods can achieve significantly superior performance than current open-source methods. Our code is available at https://github.com/emiyaning/RGCNet.
I Introduction
With the advent of consumer-grade omnidirectional cameras such as the Ricoh Theta, Samsung Gear 360, and Insta360, the acquisition of panoramic images has been significantly simplified. The $180^{\circ}\times 360^{\circ}$ field of view offered by panoramas renders them particularly valuable for indoor 3D perception applications[1]. To enable 3D understanding of indoor environments from a single omnidirectional image, depth estimation becomes a fundamental requirement[2, 3, 4, 5, 6, 7, 8]. This potential for inferring entire scene structure from a single panorama has motivated active research on panoramic depth estimation.
Figure 1: 3D visualization of panorama depth estimator’s predictions on the Stanford2d3d dataset. The left side shows the ground-truth visualization, middle column shows the visualization of Panoformer’s prediction, right side shows the visualization of our framework’s prediction.
However, the ultra-wide field of view in panoramic imaging inherently introduces object distortion. To address this issue, current panoramic depth estimation approaches primarily employ two kinds of pathway[9]. One strategy designs dedicated feature extractors motivated by panoramic imaging principles to handle distortion directly[10, 11]. The other strategy projects the equirectangular panorama (ERP) into six cubemap faces, employs feature encoders separately on both the ERP and cubemap projections [12], fuses their features, and subsequently predicts depth. These existing methods have achieved remarkable achievements in the field of depth estimation of panoramic images. However, current panoramic depth estimation methods excessively prioritize local pixel accuracy over geometric room structures, leading to two issues: (1) Noisy ground truth data induces local overfitting that propagates errors to adjacent regions; (2) Local correlations cause inaccurate predictions at 3D discontinuity regions (e.g., wall corners). As described in Fig. 1, the pre-trained panorama depth estimator PanoFormer[10] shows limited performance on the wall corners while easily be disturbed by noisy ground truth data. These two problems caused by ignoring room geometry seriously hinder the application of existing depth estimation models to real-world indoor scenarios.
To address existing problems in panoramic depth estimation, this paper proposes a depth estimation framework named as room geometry constrained depth estimation network (RGCNet). The proposed RGCNet based on multi-task learning unifying three complementary tasks: depth estimation, room layout estimation, and background segmentation. This architecture employs a shared panoramic encoder to extract multi-scale features, with three dedicated decoders generating: coarse depth map, room layout map, and background segmentation mask from these features. The depth estimation branch retains the architecture of existing high-performance depth estimators [10, 11]. However, our RGCNet treats the output of the depth decoder as a first-stage prediction and refines it using the room geometry constrain which generated from room layout and background segmentation. In our framework, the initial depth estimate and room layout are jointly utilized to compute the background depth maps. By incorporating geometric constraints of the room, these derived background depth maps are not only more robust to noise in the ground-truth depth but also yield higher accuracy in structured regions such as walls and corners. Furthermore, RGCNet leverages an end-to-end multi-task framework to decode the background depth and fuse it with the coarse depth prediction in a single efficient step. To generate the background map, a proposed room-geometry-based background depth resolving strategy leverages the layout and depth decoder’s output to calculating the corresponding background depth. To fuse the background and depth decoder’s prediction, a background-segmentation-guided fusion strategy derives fusion weights for the background and coarse depth maps from the segmentation decoder’s predictions. In general, our contributions can be summarized into the following three points:
  • A room-layout constrained depth estimation framework RGCNet has been proposed to use room structural geometry to get corrected depth predictions.
  • A background depth map resolved strategy has been proposed, which extract the room structural geometry information from multi-task decoder’s prediction.
  • A adaptively fusion strategy has been proposed, which adaptively based on extracted the geometry information to refine the depth prediction.
II Related Works
A core challenge in panoramic depth estimation is image distortion. Existing research primarily employs three approaches: 1) estimating depth solely on a single projection; 2) projecting the panorama onto multiple modalities for depth estimation; and 3) using generated background depth to guide estimation.
II-A Single Projection Inputs
Panoramic images employ spherical representations with $180^{\circ}$ vertical and $360^{\circ}$ horizontal fields of view. Processing typically involves projecting these spherical images onto 2D planes through perspective mapping. Common modalities include equirectangular[13, 14, 10, 11, 15, 16, 17], tangent[18, 19, 20], and icosahedron projections[21, 22].
For depth estimation, equirectangular projection predominates. Omnidepth[13] introduced RectNet for efficient equirectangular feature extraction and depth prediction. ODE-CNN[14] proposed a hardware-software co-design system comprising: 1) a panoramic camera with binocular depth sensors, and 2) a depth estimation model leveraging RectNet baseline enhanced with spherical feature transformers between encoder-decoder stages and deformable convolutional spatial propagation for final prediction. Following vision transformers’ emergence, specialized models have proliferated: Panoformer[10] developed reference-point window self-attention for equirectangular features within an encoder-decoder architecture; Egformer[11], SGFormer[15], GLPanoDepth[23] subsequently incorporated global receptive fields and spherical geometry constraints. To address panoramic depth data scarcity, self-supervised methods [16, 17] operate directly on equirectangular RGB images.
Alternatively, tangent-view approaches [18, 19, 20] project panoramas onto multiple views, estimate per-view depth, and spatially composite patches into panoramic depth maps. Vertically compressed methods [24, 25, 26] adapt layout estimation techniques, reducing panoramas to 1-pixel height and employing Bi-LSTMs or self-attention for depth estimation.
II-B Bi-Projection Inputs
Methods using bi-projection inputs project panorama images onto two distinct perspectives. Of the bi-projection input methods, equirectangular projection is usually used as the primary perspective by default. These methods usually use shared or dedicated branch networks to predict corresponding depth maps. Then, the another perspective-specific depth estimate are reprojected to the equirectangular domain and composited with directly predicted equirectangular depth maps to yield refined depth estimations. Representative examples include BiFuse[12], BiFuse++[27], and UniFuse[28], which project onto cube maps; HRDFuse[29] and GA360Fuse[30], which fuse equirectangular and tangent depth predictions; and Elite360D[31], which fuses equirectangular and ICOSAP[22] perspectives for enhanced accuracy. Intuitively, such methods seem better suited to handle distortion than single-view panoramic depth estimation. However, current public dataset rankings show that supervised learning on pure equirectangular images achieves superior performance. Moreover, dual-view projection inevitably introduces significant additional computational overhead during training and inference. Therefore, the benefits of this approach may not justify its computational cost.
II-C Background Based Methods
Current state-of-the-art methods demonstrate strong quantitative performance across benchmarks. However, their 3D depth visualizations frequently exhibit structural inaccuracies in room geometry and over-smoothed corners. To address this limitation, recent approaches[32] integrate room structural priors into panoramic depth estimation. The core challenge lies in efficiently computing accurate background depth maps, which require precise room layout estimation.
Indoor panoramic layout estimation aims to detect room wall boundaries from input panoramas. Existing research predominantly adopts the Manhattan World assumption. LayoutNet[33] directly predicts per-pixel corner and boundary probability maps. Conversely, Dula-Net[34] decodes panoramas into equirectangular and perspective views, extracting wall-floor/ceiling boundaries via semantic segmentation. EquiConv[35] employs specialized convolutions to generate corner/edge probability maps. These early methods share a 2D segmentation paradigm, outputting full-resolution probability maps—an inefficient approach given layout estimation only requires boundary/corner identification. HorizonNet[36] thus reformulates the task as 1D sequence prediction using Bi-LSTMs, while HoHoNet[29] employs multi-head self-attention for similar 1D representation. LED-Net[37] further predicts layouts from horizontally compressed representations. Subsequently, AtlantaNet[38] projects panoramas onto ceiling/floor planes to avoid occlusion, predicting layouts via amalgamated projections. DMH-Net[39] extends this by mapping to cubemap faces before predicting boundaries and corners.
A critical secondary challenge involves fusing background depth with panoramic depth predictions. The recent BGDNet[32] incorporates room geometry as background depth when predicting the final depth map. However, it needs pretrained HorizonNet and SAM[40] to extract the background depth map from input panorama image. In this paper, we adopt a multi-task learning approach to integrate room layout and segmentation predictions needed for background depth calculation within a single framework. We also design a background depth calculation strategy based on background segmentation and a fusion strategy for coarse and background depth guided by background segmentation weights. Our method is detailed in the following sections.
Figure 2: The structure diagram of our proposed room geometry guided depth estimation framework. In terms of model structure, our framework includes a shared panorama encoder and three task-corresponding decoders. Based on the obtained layout map, coarse depth map, and background-segmentation map, the proposed framework decode fine-grained depth prediction.
III Methodology
The room structural regularized depth estimation framework proposed in this paper is shown in Fig.2. From the view of neural network architecture, the framework includes a shared panorama encoder and three task-corresponding decoder modules. Moreover, our framework contain a room geometry-based background depth resolving strategy and a background-segmentation-guided fusion mechanism. The panorama encoder extract the multi-stage feature sets $F=\{f^{2^{i+1}\times}_{i}|i=1,2,3,4\}$ from the panorama image. The decoder modules corresponding to the subsequent three tasks use these features as input to estimate the corresponding results. The proposed room geometry-based background depth resolving strategy calculate the background depth based on panorama depth, room layout, background segmentation predictions. Then, background-segmentation-guided fusion mechanism regularize the depth decoder’s prediction based on background depth map and background segmentation predictions. In the following subsections we will introduce each module in detail.
III-A Architecture
Figure 3: The structure diagram of layout feature aggregation module and pixel feature aggregation module.
Panorama Encoder: In this paper, the framework we proposed uses the backbone proposed by PanoFormer as a common feature encoder. The backbone consists of the panorama transformer block and convolution 2D layer based downsample layer. The panorama transformer block contains a window self-attention mechanism designed for the panorama imaging process and a feed-forward layer designed based on depthwise separable convolution[41]. The entire backbone consists of an input projection layer and $4$ stage blocks. The input projection layer consists of a convolution layer whose kernel size is $3\times 3$ with stride of $2$, a 2D batchnorm layer and a ReLU activation function. The subsequent $4$ stage blocks all consist of a panorama transformer block and a double-downsampling layer composed of a 2D convolutional network. Each stage block will output the corresponding double-downsampled feature map result.
Layout Decoder: Considering that layout estimation is a non-pixel-level task while depth estimation and background segmentation are both pixel-level tasks, and some existing work [42] also mentioned that different tasks may have different requirements for features of different scales. To this end, we design a specific layout FA module for the layout task to further process the features extracted by the panorama encoder. The structure diagram of the layout feature aggregation module for the layout task is shown in Fig.3 (a) . This module use the height compress convolution with spatial stride $(2,1)$ to compress the feature maps $f^{4\times}_{1},f^{8\times}_{2},f^{16\times}_{3},f^{32\times}_{4}$. Then the compressed feature map is flattened to obtain a feature vectors set $\hat{f}_{1},\hat{f}_{2},\hat{f}_{3},\hat{f}_{4}$. We then cat these flattened features together and, following the Hohonet[25] model design experience, use a multi-head self-attention module to extract global feature information from the flattened feature vectors and output the final feature vector $\hat{f}_{cat}$. Based on the output features of the layout FA module, the layout decoder uses boundary and corner heads, consisting of a convolution 1D layer, a ReLU layer, and a batchnorm 1D layer, to predict the final room layout $S_{room}$.
Depth Decoder: Considering that both depth estimation and background segmentation tasks are pixel-level tasks, we follow the design idea of Hohonet and let them use a common pixel-level feature aggregation to process the multi-stage panorama feature. The structure of the pixel-level feature aggregation module we use is shown in Fig. 3 (b). This module consists of a panorama transformer block and an upsampling module. Given multi-scale feature map set $f^{1}_{4\times},f^{2}_{8\times},f^{3}_{16\times},f^{4}_{32\times}$, this module’s feature integration process can be described as:
$\hat{f}=\prod^{3}_{i=1}[UP(PT(f^{2^{i}\times}_{i}))+f^{2^{i+1}\times}_{i+1}]$ (1)
On the integrated multi-scale feature $\hat{f}$, the depth estimation decoder predict the coarse depth maps $S^{p}_{depth}=\{d_{ij}|i=1,2,...,W;j=1,2,...,H\}$.
background-Segmentation Decoder: As mentioned above, the background segmentation task and the depth estimation task share a pixel-level feature aggregation module. The structure diagram of the pixel-level feature aggregation module is shown in Fig. 3(b). Therefore, in our framework, the background-segmentation decoder itself has only one prediction head consisting of a convolution 2D layer, a ReLU function, and a batchnorm 2D layer. This prediction head predicts the corresponding 2D background segmentation result $S_{seg}=\{0\leq p_{ij}\leq 1|i=1,2,...,W;j=1,2,...,H\}$ based on the integrated multi-scale feature $\hat{f}$.
III-B Room Geometry-based Background Depth Resolving Strategy
Figure 4: P is the camera center, A and B are the upper and lower boundary points of the wall corresponding to point P, and D is an arbitrary point on the wall plane. The lengths of AB in the image are known, and the corresponding angles $\phi_{c}$ and $\phi_{f}$ can be calculated based on spherical camera geometry. Based on the depth of point P predicted by the depth decoder, $d_{c}$ and $d_{f}$ can be calculated.
Given a room layout estimation prediction $S_{room}\in R^{W\times 3}$ to resolve the corresponding background depth, we need to obtain the camera-to-ground distance information. Existing methods often assume a fixed camera height when calculating background depth based on layout information. This assumption is stable and reliable in virtual rendering datasets, but it is not necessarily reliable when collecting data in the real world using a panoramic depth camera. To obtain more stable and reliable camera height information for each scene, we first designed a camera height resolution strategy.
As shown in the Fig. 4, define the coordinates of the camera origin $O$, and draw a normal line from the camera plane to the image plane, which intersects the wall plane at point $P$. Then draw a straight line perpendicular to the camera plane at point $P$, which intersects with the ceiling-wall boundary line in the room layout prediction at point $A$, and intersects with the floor-wall boundary line at point $B$. What our camera height calculation strategy needs to do is to calculate the camera height in the Fig. 4. Our method use the predictions of the depth decoder and room layout decoder as the input of camera height calculating method. The room layout predictions are responsible for providing the image plane coordinates of the intersection points $A$ and $B$ in the panorama image and depth maps. Based on the results of the depth maps prediction, we can get the distance between the corresponding line segments $OA$ and $OB$. On this basis, the distance from the camera to the ceiling and the camera height can be conveniently calculated using the following formula:
$|AP|=d_{c}\times sin(\phi_{c})$ (2)
$|PB|=d_{f}\times sin{\phi_{f}}$
The corresponding angles $\phi_{c}$ and $\phi_{f}$ can be calculated using the geometric imaging principle of a panoramic camera. Assuming that the vertical coordinate of point $A$ is $u_{ceil}$ and the vertical coordinate of point $B$ is $u_{floor}$, the calculation formulas for the two angles can be obtained using the following formulas:
[…]
After determining the camera height and the distance between the camera and the ceiling, we can use the wall information from the room layout predictions to calculate the corresponding background depth maps for each pixel belonging to the background wall. The depth maps for the ceiling and floor can be simply calculated based on the angles calculated from the pixels using the following formula:
$d^{i}_{floor}=\frac{|PB|\times H}{(u^{i}_{floor}-0.5H)\times\pi}$ (4)
$d^{i}_{ceil}=\frac{|AP|\times H}{(0.5H-u^{i}_{ceil})\times\pi}$
For any point $D$ on the wall plane that is tilted toward the floor, assume its equirectangular plane coordinates on the panorama image are $(u_{i},v_{i})$. First, we can calculate the horizontal and vertical offset angles $\rho_{D}$ and $\phi_{D}$ of this point relative to the camera center $P$ using the following formulas:
[…]
$\rho_{D}=(1-\frac{v_{i}}{W})\times\pi$
Based on the two deflection angles, to obtain the depth of $D$, $d_{D}=|OD|$, we first draw a perpendicular line through point $D$ to line segment $AB$ and intersect $AB$ at point $E$. First, based on the vertical offset angle $\phi_{D}$, we can obtain $|OE|=cos(\phi_{D})|OP|$, where the value of $|OP|$ can be directly obtained from depth prediction. Then, based on the horizontal offset angle $\rho_{D}$ and $|OE|$, we can simply solve $d_{D}=cos(\rho_{D})|OE|$. Similarly, any point on the wall plane that is deflected toward the ceiling can also use a similar method to solve for its corresponding depth value. Finally, combining all the solved depth values, we can obtain $S_{back}$.
III-C Background-segmentation-guided Fusion Mechanism
As mentioned above, based on the corase depth map and the background depth map, our framework needs to fuse the predicted depth $S^{p}_{depth}$ and the background depth $S_{back}$ to obtain an accurate final depth estimation result $S^{final}_{depth}$. Considering that $S_{back}$ is naturally suitable as an upper bound for depth estimation when the layout prediction is relatively accurate. Then $S_{back}$ as an upper bound can be used to ensure that the depth of pixels originally belonging to the rescue, ceiling, ground and other areas do not exceed the area of the room itself. The depth value of the pixel belonging to the foreground object area can be as close to $S^{p}_{depth}$ as possible. Therefore, in the process of fusing the two depth maps, we need a weight to determine whether the current pixel is an object belonging to the foreground area or the background area. The probability map $S_{seg}$ predicted by the background-segmentation task in our framework can just be used as such a fusion weight. Therefore, the entire fusion process can be described as the following formula:
$S^{final}_{depth}$ $=\{d^{final}_{ij}|i=1,2,.,W;j=1,2,.,H\};$ (6)
$etl.$ $\ d^{final}_{ij}=d^{back}_{ij}\times p_{ij}+d^{p}_{ij}\times(1-p_{ij});$
$p_{ij}\in S_{seg};$
$d^{back}_{ij}\in S_{back};$
$d^{p}_{ij}\in S^{p}_{depth};$
III-D Objective Function
[…]
When training the layout branch, we refer to the settings of HorizonNet and use the binary cross entropy and L1 loss function to calculate. When training the depth estimation branch, we used the same Huber[43] (or Berhu[44]) loss and gradient loss[45] as the baseline Panoformer. The hyper-parameters within the objective functions of these two branches were also consistent with those of HorizonNet and PanoFormer.
Considering that existing panoramic image datasets usually do not contain background segmentation annotations. When training the background segmentation decoder, we first use the background depth map $S_{back}$ predicted by the framework and the corresponding ground-truth depth map $S^{gt}_{depth}$ to calculate the corresponding background segmentation label $S^{gt}_{seg}$. The specific solution formula is as follows:
$D^{res}=|S^{gt}_{depth}-S_{back}|;$ (7)
$S^{gt}_{seg}=\{p^{gt}_{ij}=\Gamma(d^{res}_{ij}<\gamma)|d^{res}_{ij}\in D^{res}\}$
where $\Gamma$ is a binary function, which takes the value 1 when the condition in the brackets is true and takes the value 0 when it is false.
When optimizing the partial network for the background-segmentation task, we use a simple focal loss as our object function, which is calculated as follows:
$L_{seg}=-\frac{1}{M}\sum^{M}_{u=1}\frac{1}{H*W}\sum^{W}_{i=0}\sum^{H}_{j=0}\alpha(1-\hat{p}_{ij})^{\eta}log(\hat{p}_{ij});$ (8)
where $\hat{p}_{ij}=p_{ij}$ if $p^{gt}_{ij}=1$ else $\hat{p}_{ij}=1-p_{ij}$. Following the common setting of focal loss, we set the $\alpha$, $\beta$ as 0.5, 2.0 relatively.
In summary, the overall object function of our proposed framework can be described as:
$L_{all}=\lambda_{1}*L_{layout}+\lambda_{2}*L_{depth}+\lambda_{3}*L_{seg}$ (9)
where $\lambda_{1},\lambda_{2},\lambda_{3}$ are three manual setting parameters.
IV Experiment
IV-A Experiment Setting
In this section, we introduce the experiments to verify the proposed depth estimation framework. In our experiments, we selected three datasets: Stanford2D-3D[46], MatterPort3D[47], and Structure3D[48]. Stanford2D-3D and MatterPort3D are two datasets collected in the real world, while the Structure3D dataset is a synthesized dataset. Real-world datasets contain a large amount of noise points introduced during the acquisition process as described in OnimiDepth[13]. This noise introduces local jumps that severely hinder model performance. Therefore, before training models on Stanford2D3D and Matterport3D, we used a proposed layout-based dataset denoise strategy to constrain the depth of these noise points to within the room.
Stanford2D-3D: The Stanford2D-3D dataset contains 1413 panoramic images collected from 3 types of buildings divided into 6 large areas. We follow the official practice to divide the dataset into training and test sets, and downsample all depth maps and RGB images to $512\times 1024$ size images.
Matterport3D: Matterport3D contains 10,800 panoramic images collected from 90 different rooms. The camera used in the collection process of this dataset is Matterport’s Pro 3D camera. In this part of the dataset, we also use 61 room images for training and 29 room images for testing. All RGB and depth images are also downsampled to $512\times 1024$ size.
Structured3D: The Structured3D dataset contains 196K rendered panoramic images and corresponding depth maps, covering 12,835 rooms in 3,500 scenes. Each room is created manually using CAD models of furniture, which are in real-world dimensions and used in real production. In this dataset, we follow the official recommended setting [48], using the data of the first 3000 scenes as the training set, the data of 3000-3250 scenes as the validation set, and the data of the last 250 scenes as the test set.
Metric: Following the previous works, we use some standard evaluation metrics, which include: relative error (abs rel), squared relative error (sq rel), root mean squared error (rmse), and three threshold percentage $\delta<\varsigma^{t}(\varsigma=1.25,t=1,2,3)$ denoted as $\delta_{t}$.
Dataset denoise strategy: Considering that datasets collected in the real world contain a large amount of noisy points, we have also designed a denoising strategy for panoramic depth map datasets. As mentioned earlier, our method uses the outputs of the layout decoder and depth decoder to calculate the corresponding background depth map. Here, we use this calculated background depth map to constrain the dataset collected in the real world. First, for areas in the real-world depth ground-truth maps where measurement fails, we directly replace them with the background depth value. Then, our method sets a threshold to judge the ground-truth depth. If the depth point in the ground-truth dataset of the original data is converted to 3D space and its coordinates are outside the room and the distance from the room wall is more than 1 meter, we identify the pixel as a noise pixel. For these pixels, we also use the depth in the background depth map as a replacement.
Training setting: We use Adam as the optimizer, and the parameters of the optimizer are basically the basic settings of the pytorch framework. For the learning rate scheduling strategy, we choose one-cycle[49], set the initial learning rate to 0.0001, and the minimum learning rate to 0.0000001. Our hardware experimental platform is configured with AMD Epyc 7003 CPU and 4-card RTX 4090 GPU. During the training process, we set the batch size on each card to 2. In the data set enhancement part, we used random horizontal angle rotation and random horizontal flipping with reference to Panoformer[10]. For horizontal angle rotation, we set the interval of random angle to $\frac{\pi}{4},\frac{\pi}{4}]$.
Parameter setting: The training objective function used by our framework is shown in Section 4.4, which contains three subcomponents: $L_{front}$, $L_{depth}$, and $L_{layout}$. Considering that some datasets do not contain complete layout annotations, and our framework is extremely dependent on the accuracy of layout estimation, we will first extract all layout annotations from the three selected datasets to form a large layout dataset to pretrain the layout task. Then, when training the depth estimation task on specific datasets, we decide whether to finetune the weights of the layout and feature encoder parts based on whether the dataset contains complete layout annotations. Specifically, the Stanford2D-3D data does not provide complete layout annotations, so we set the corresponding hyper-parameter $\lambda_{1}$ to 0 when training our framework, and the corresponding $\lambda_{2},\lambda_{3}$ to 1.0 and 0.4. The Matterport3D and Structure3D datasets provide complete layout annotations, so we set $\lambda_{1},\lambda_{2},\lambda_{3}$ to 0.01, 1.0, and 0.4, respectively. In addition, some hand-tuned hyper-parameter settings related to this study, such as $\rho_{1}$ and $\rho_{2}$ in equation was set to 1.0 and 0.5. Other hyper-parameters that are not relevant to this study are set according to the common specifications in the current field.
IV-B Performance Comparison
Figure 5: The 3D visualize results of final depth estimation. For each scene, we selected three perspectives: top view, side view, and internal perspective to display the three-dimensional visualization effect of the point cloud converted from the depth map.
IV-B1 Quantification Comparison with SoTA Depth Estimation Methods
TABLE I: Quantification comparison with State of the Art depth depth estimation methods on Matterport3D
Dataset Method Pub’ Year Classic Metrics
$\delta_{1}$ $\delta_{2}$ $\delta_{3}$ RMSE MRE MAE
Matterport3D EGFormer[11] ICCV 2023 0.8158 0.9390 0.9735 0.6025 0.1517 0.1473
OminiFusion[19] CVPR 2022 0.9040 0.9757 0.9919 0.4261 0.0552 0.0900
Bifuse[12] CVPR 2020 0.8452 0.9319 0.9632 0.6295 0.2408 0.3470
UniFuse[28] IEEE RAL 2021 0.8897 0.9623 0.9831 0.4941 - 0.2814
HRDFuse[29] CVPR 2023 0.9162 0.9669 0.9844 0.4433 0.0936 0.0967
RSDNet[50] CVPR 2022 0.443 0.097 0.248 0.906 0.971 -
SN360[51] IEEE Access 2025 0.4483 - - 0.9392 0.9808 0.9932
GLPanoDepth[23] IEEE TIP 2024 0.8641 0.9561 0.9808 0.5223 - 0.2998
Panoformer[10] ECCV 2022 0.9184 0.9804 0.9916 0.3635 0.0571 0.1013
SGFormer[15] IEEE TCSVT 2025 0.8946 0.9642 0.9859 0.4790 - 0.2748
Ours 0.9199 0.9820 0.9984 0.2436 0.0425 0.0829
TABLE II: Quantification comparison with State of The Art depth estimation methods on Stanfor2d3d
Dataset Method Pub’ Year Classic Metrics
$\delta_{1}$ $\delta_{2}$ $\delta_{3}$ RMSE MRE MAE
Stanford2d3d EGFormer[11] ICCV 2023 0.8185 0.9338 0.9736 0.4974 0.1408 0.1528
OminiFusion[19] CVPR 2022 0.8940 0.9714 0.9900 0.3715 0.0543 0.0961
Bifuse[12] CVPR 2020 0.8660 0.9580 0.9860 0.4142 0.1209 0.2343
UniFuse[28] IEEE RAL 2021 0.8711 0.9664 0.9882 0.3691 - 0.2082
HRDFuse[29] CVPR 2023 0.8941 0.9778 0.9923 0.3452 0.0503 0.0984
FreDSNet[52] 2022 0.8424 0.9583 0.9863 0.2727 0.0952 0.1327
RSDNet[50] CVPR 2022 0.394 0.098 0.209 0.903 0.974 -
SN360[51] IEEE Access 2025 0.2917 - - 0.9369 0.9846 0.9942
GLPanoDepth[23] IEEE TIP 2024 0.9015 0.9793 0.9901 0.3493 - 0.1932
Panoformer[10] ECCV 2022 0.9394 0.9838 0.9941 0.3083 0.0405 0.0619
SGFormer[15] IEEE TCSVT 2025 0.8998 0.9693 0.9908 0.3406 - 0.2017
Ours 0.9479 0.9857 0.9943 0.2359 0.0285 0.058
TABLE III: Quantification comparison with State of the Art depth estimation methods on Structured3D
Dataset Method Pub’ Year Classic Metrics
$\delta_{1}$ $\delta_{2}$ $\delta_{3}$ RMSE MRE MAE
Structure3d EGFormer[11] ICCV 2023 0.7979 0.9071 0.9455 0.6841 0.4509 0.2205
OminiFusion[19] CVPR 2022 0.6921 0.8831 0.9501 0.4951 - 0.2981
Bifuse[12] CVPR 2020 0.8594 0.9400 0.9672 0.5213 0.2455 0.1573
UniFuse[28] IEEE RAL 2021 0.8542 0.9399 0.9676 0.5016 0.2319 0.1506
HRDFuse[29] CVPR 2023 0.7561 0.9161 0.9631 0.4061 - 0.2451
BGDNet[32] CVPR 2024 0.8336 0.9377 0.9731 0.3490 - 0.1656
SGFormer[15] IEEE TCSVT 2025 0.9613 0.9896 0.9957 0.2429 - -
PanoFormer[10] ECCV 2022 0.8943 0.9536 0.97431 0.3017 0.1201 0.1546
Ours 0.9679 0.9907 0.9983 0.1935 0.0414 0.0613
Figure 6: Visualization of background depth map guided dataset denoising effect diagram.
[…]
The Stanford2d3d dataset is a relatively early dataset, and various existing depth estimation methods have achieved relatively good performance on this dataset. However, on this dataset, the noise processing strategy and multi-task depth estimation framework proposed in this paper still show excellent effectiveness. Specifically, our method significantly outperforms the current SOTA method by 3.68 percentage points in the RMSE, the evaluation indicator that the industry attaches the most importance to.
At the same time, in MRE and MAE, which are both numerical indicators, the performance advantage of our method over the existing methods is not as obvious as RMSE, but it is still intuitive, which further proves the robustness and superiority of the method proposed in this paper. In the three indicators $\delta_{1},\delta_{2}$ and $\delta_{3}$ at the ratio level, our method also has a leading advantage, but this advantage is not as obvious as the numerical indicators. We believe that this is because the existing methods have reached a very high level in the three indicators at the ratio level, so it is more difficult to achieve higher indicators.
On the Matterport3D dataset, our method also achieved the current best performance in RMSE, the core evaluation indicator of depth estimation, which fully demonstrated the effectiveness of the proposed method and data processing strategy in key accuracy metrics. However, on the relative error threshold indicator ($\delta<$ 1.25), our method failed to achieve the best results. We believe this is related to the introduced layout constraints: this constraint tends to optimize the absolute error between the prediction and the true value, but in areas with small true depth values, this optimization may result in a limited range of change in the predicted value, making it more difficult to meet the strict 1.25 times relative error requirement. Despite this, we still achieved a level close to SOTA on this indicator. Importantly, this method is significantly ahead of existing work in both the core rmse and rmslog indicators.
On the Structured3d dataset, we mainly compare with BGDNet, which uses a similar idea. The results released in the BGDNet paper are trained on the replica and then val on the structure3d dataset, while our method is trained on Structured3d and then validated on the corresponding validation set. This comparison may not be fair, but considering that the BGDNet paper is not open source, we can only use this method for comparison. Due to the differences in the settings of train and val, the method proposed in this paper has a very obvious advantage in the six indicators of concern.
TABLE IV: Quatitative analysis of our proposed dataset denoise method.
Dataset Setting Methods RMSE MRE MAE
Without Denoise PanoFormer 0.3083 0.0405 0.1013
Our Methods 0.2673 0.0375 0.0912
Denoised Dataset PanoFormer 0.2872 0.0392 0.0892
Our Methods 0.2359 0.0285 0.0580
TABLE V: The training impact of each task decoder to depth decoder
Depth
Decoder
Front-Seg
Decoder
Layout
Decoder
RMSE MRE MAE
0.3635 0.0571 0.1013
0.3621 0.0531 0.1011
0.3656 0.0582 0.1101
0.3592 0.0511 0.1002
TABLE VI: The accuracy of resolved background depth maps.
Methods RMSE MRE MAE
PanoFormer’s Background Region 0.3214 0.0623 0.1123
Decode from layout prediction 0.0043 0.0032 0.0023
Decode from layout ground-truth 0.0001 0.0001 0.0001
TABLE VII: The comparison of our proposed fusion methods and BGDNet
Methods RMSE MRE MAE
Baseline 0.3635 0.0571 0.1013
Fusion BGDNet 0.2823 0.0498 0.0922
Fusion ours 0.2436 0.0425 0.0829
Figure 7: The experiment results of explore the inference of hyper-parameter $\lambda_{2}$ and $\lambda_{3}$. The x-axis in this figure represents the value of the corresponding hyper-parameter setting, while y-axis represents the RMSE value of our proposed framework’s depth prediction.
IV-B2 Visualize Comparison with SoTA Depth Estimation Method
In this section, we perform visualization operations on the scene as shown in the figure. In this set of experiments, we selected the baseline panoformer in this paper as the comparison object. It is worth noting that in this set of experiments, panoformer and the depth estimation framework proposed in this paper both use denoised datasets during training. In this set of experiments, we selected several representative scenes on the validation set for demonstration. From the corresponding 3D visualization results, it can be seen that the depth estimated by our method is significantly more accurate in corners and walls. At the same time, thanks to the prior knowledge of room layout introduced by us, the 3D visualization of the prediction results of our method is more reasonable in terms of overall structure. At the same time, due to the existence of our dataset denoising strategy, panoformer and our method do not have the phenomenon of overfitting of noise points in patches as shown in Fig. 1. The existence of these phenomena fully demonstrates the effectiveness of the depth estimation method we proposed.
IV-C Visualize and Quatitative analysis of Dataset Denoise Method
IV-C1 Visualize Analysis of Denoise Method
We conducted a visualization experiment as shown above for the panoramic depth images of the 2D-3D-S dataset collected in the real world. After mapping each pixel in the panoramic depth images to 3D space, we found that there are a large number of noise points in each scene. For these noise points, the processing strategy of the existing panoramic depth estimation method is to directly preset a maximum distance threshold. In the 2D-3D-S dataset, this threshold is set to 10 meters. The depth values of all areas greater than this threshold will be artificially set to 0. On the one hand, this processing method destroys the correlation between the pixels of the original panoramic depth map, thereby affecting the learning process of the model. On the other hand, it will also cause some areas in the scene with depth values less than 10 but should fall inside the room to fall outside the room. In order to preserve the pixel correlation of the processed panoramic depth map and prevent the area falling inside the room from falling outside the room, we proposed a strategy to preprocess the indoor panoramic depth map using pretrained layout. In this strategy, we first collected as many parts of the existing dataset that contain layout annotations as possible to form a layout dataset of a relatively reasonable size to pretrain the layout estimation part of the framework proposed in this paper. Considering that the layout prediction branch can already obtain a high 3D IoU in our experiments, we directly use this part of the prediction results to process the depth estimation dataset in the actual processing process. In our proposed processing strategy, we first use the layout prediction branch to estimate the layout of the current room, and then use the panorama depth ground-truth to count the height of each wall of the room, and use this height result to solve the 2D layout into the corresponding background depth map. Finally, we use this background depth map to constrain the original panorama depth ground-truth as shown in Fig. 6. In this constrained process, we use the background ground-truth as a mask to ensure that the depth of the background region does not exceed the range of the room itself.
IV-C2 Quatitative Analysis of Denoise Method
We use the Stanfor2d3d dataset as the object to verify the effectiveness of the proposed noisy image processing strategy. In the experiment, we use our proposed framework and PanoFormer to perform training and val on the processed datasets before and after processing. The experimental results are shown in Table IV. From the results in Table IV, the performance of PanoFormer and our proposed framework trained on the processed dataset is significantly higher than that trained on the original dataset, which fully demonstrates the performance improvement brought by our data processing strategy.
IV-D Ablation Study
IV-D1 The Effect of Framework’s Each Component
Figure 8: The visualize of Background segmentation. The left column of the figure is input panorama RGB images, the middle column is background segmentation predictions, the right column is corresponding ground-truth.
Figure 9: The visualize of room layout estimation. The first column shows the input RGB image. The second column shows the room layout estimation results displayed on a 2D equirectangular image. The top row in each 2D equirectangular sub-image represents the probability of a corner point in that column. The third and fourth columns show the corresponding 3D bird’s-eye views of the prediction and layout ground-truth, respectively.
In this section, we analyze the impact of each proposed component on the whole. Considering that we proposed a framework for auxiliary-task assisted depth estimation tasks, we mainly discuss two aspects of the impact of each component: 1) The impact of auxiliary-task decoder for depth estimation decoder; 2) The effect of our proposed background depth resolving and fusion strategy. The experimental dataset we chose is Matterport3D because it contains complete layout estimation annotations.
Regarding the aspect 1), the corresponding experimental results have been put into Table V. It is worth noting that the loss weight we used in this group of experiments is the combination verified in subsequent section. From the results, it can be seen that the auxiliary-task decoder has very little impact on the depth decoder during training. This is primarily because we intentionally set a low loss weight for the auxiliary-task decoder. This design stems from two considerations: first, optimizing each task in a multi-task learning framework requires extensive experimentation and manual tuning. Second, and more importantly, our framework’s primary objective is to use the auxiliary task’s output to constrain the depth decoder directly, rather than to improve it indirectly by learning more generalized features in the shared encoder.
Regarding the aspect 2) we want to discuss, we conducted corresponding experiments and put them in Table VI and Table VII. The experimental results of the background depth resolving strategy are shown in Table VI, and the experimental results of the fusion strategy used are shown in Table VII. When evaluating the background depth resolving strategy, we used the depth ground-truth provided by the Matterport3D dataset as a benchmark. When evaluating the resolved background depth, we used the ground-truth of the background segmentation as a mask to ensure that only the background region is evaluated during the eval process. Judging from the results in Table VI, our background depth resolving strategy is more accurate than the deep learning model. In terms of fusion strategy, we chose BGDNet[32] as the comparison object. In this set of experiments, the configuration of the fusion strategy used by BGDNet is consistent with the original paper. The experimental results demonstrate that incorporating background depth as a constraint effectively enhances depth estimation performance. Moreover, our proposed fusion strategy yields even more substantial improvements compared to that of BGDNet.
In summary, our discussion validates both the proposed framework and the collaborative design of its subtasks.
IV-D2 Analysis of loss weights
The impact of different loss weights hyper-parameters of the proposed framework has been shown in Fig. 7. Here, we use the MatterPort3D to conduct experiments. As the layout decoder in our framework has been pretrained, we simply fix its $\lambda_{1}$ to 0.01. For $\lambda_{2}$ and $\lambda_{3}$, we first fix the $\lambda_{3}$ to 1.0 and search the best setting of $\lambda_{2}$. Then, fix the $\lambda_{2}$ to the searched results and search the setting of $\lambda_{3}$. The Fig. 6. show that on the MatterPort3D dataset, when we set $\lambda_{2}=1.0$ and $\lambda_{3}=0.4$, the proposed framework can achieve the best performance.
IV-E Visualize Analysis of Background Segmentation and Layout Estimation
In this section, we visualize the prediction results of the front-segmentation and layout estimation task. We used the test set data of Structured3D to perform correspond experiment. The specific visualization results are shown in Fig. 8 and Fig. 9.
Fig. 8 shows the visualization results of the background segmentation task. In our framework, the prediction results of the background segmentation task are used as weight to fuse the coarse prediction and background depth. When visualizing the results of background segmentation, we mapping the prediction score maps from $[0,1]$ to $[0,255]$ and save them as gray images. Judging from the score maps shown in Fig. 8, our background branch can accurately distinguish between the background and background areas in the input image. However, in some areas that are very close to the wall and have a similar depth to the background area, the discrimination ability of our background branch still needs to be improved.
Fig. 9 shows the estimation results of our layout estimation branch. From the 2D visualization of the prediction results, we can see that the edge areas and corner points of the wall can be identified relatively accurately. However, in some areas far away from the camera, the prediction accuracy of the corner points still needs to be improved. In addition to these issues, another point worth noting is that the current room layout estimator assumes that the walls of the room are all cubic structures. This makes our method perform worse than the normal depth estimator when estimating the depth of some rooms with curved wall structures. We consider conducting more detailed research on this issue in future work.
V Conclusions
This paper has proposed a panoramic depth estimation framework with room geometry constraints. The framework employs multi-task learning, where a shared encoder extracts features that are decoded into three outputs: coarse depth maps, background segmentation masks, and room layouts. These predictions enable our method to initially reconstruct background depth from the layout information. Subsequently, background depth and coarse depth are fused using the background segmentation mask as a weighting mechanism, ultimately generating the final depth prediction. Extensive experiments on real-world and synthetic datasets demonstrate significant performance improvements over current methods.

### v2

@@
-An End-to-End Room Geometry Constrained Depth Estimation Framework for Indoor Panorama Images
+PAGCNet: A Pose-Aware and Geometry Constrained Framework for Panoramic Depth Estimation
 Abstract
-Predicting spherical pixel depth from monocular $360^{\circ}$ indoor panoramas is critical for many vision applications.
-However, existing methods focus on pixel-level accuracy, causing oversmoothed room corners and noise sensitivity.
-In this paper, we propose a depth estimation framework based on room geometry constraints, which extracts room geometry information through layout prediction and integrates those information into the depth estimation process through background segmentation mechanism.
-At the model level, our framework comprises a shared feature encoder followed by task-specific decoders for layout estimation, depth estimation, and background segmentation.
-The shared encoder extracts multi-scale features, which are subsequently processed by individual decoders to generate initial predictions: a depth map, a room layout map, and a background segmentation map.
-Furthermore, our framework incorporates two strategies: a room geometry-based background depth resolving strategy and a background-segmentation-guided fusion mechanism.
-The proposed room-geometry-based background depth resolving strategy leverages the room layout and the depth decoder’s output to generate the corresponding background depth map.
-Then, a background-segmentation-guided fusion strategy derives fusion weights for the background and coarse depth maps from the segmentation decoder’s predictions.
-Extensive experimental results on the Stanford2D3D, Matterport3D and Structured3D datasets show that our proposed methods can achieve significantly superior performance than current open-source methods.
-Our code is available at https://github.com/emiyaning/RGCNet.
+Explicitly modeling room background depth as a geometric constraint has proven effective for panoramic depth estimation.
+However, reconstructing this background depth for regular enclosed regions in a complex indoor scene without external measurements remains an open challenge.
+To address this, we propose a pose-aware and geometry-constrained framework for panoramic depth estimation.
+Our framework first employs multiple task-specific decoders to jointly estimate room layout, camera pose, depth, and region segmentation from a input panoramic image.
+A pose-aware background depth resolving (PA-BDR) component uses tasks decoder’s prediction to resolve the camera pose.
+Subsequently, the proposed PA-BDR component uses the camera pose to compute the background depth of regular enclosed regions and uses this background depth as a strong geometric prior.
+Based on the output of the region segmentation decoder, a fusion mask generation (FMG) component produces a fusion weight map to guide where and to what extent the geometry-constrained background depth should correct the depth decoder’s prediction.
+Finally, an adaptive fusion component integrates this refined background depth with the initial depth prediction, guided by the fusion weight.
+Extensive experiments on Matterport3D, Structured3D, and Replica datasets demonstrate that our method achieves significantly superior performance compared to current open-source methods.
+Code is available at https://github.com/emiyaning/PAGCNet.
 I Introduction
-With the advent of consumer-grade omnidirectional cameras such as the Ricoh Theta, Samsung Gear 360, and Insta360, the acquisition of panoramic images has been significantly simplified.
-The $180^{\circ}\times 360^{\circ}$ field of view offered by panoramas renders them particularly valuable for indoor 3D perception applications[1].
-To enable 3D understanding of indoor environments from a single omnidirectional image, depth estimation becomes a fundamental requirement[2, 3, 4, 5, 6, 7, 8].
+To enable 3D understanding of indoor environments from a single omnidirectional image, depth estimation becomes a fundamental requirement[19].
 This potential for inferring entire scene structure from a single panorama has motivated active research on panoramic depth estimation.
-Figure 1: 3D visualization of panorama depth estimator’s predictions on the Stanford2d3d dataset.
-The left side shows the ground-truth visualization, middle column shows the visualization of Panoformer’s prediction, right side shows the visualization of our framework’s prediction.
-However, the ultra-wide field of view in panoramic imaging inherently introduces object distortion.
-To address this issue, current panoramic depth estimation approaches primarily employ two kinds of pathway[9].
-One strategy designs dedicated feature extractors motivated by panoramic imaging principles to handle distortion directly[10, 11].
-The other strategy projects the equirectangular panorama (ERP) into six cubemap faces, employs feature encoders separately on both the ERP and cubemap projections [12], fuses their features, and subsequently predicts depth.
-These existing methods have achieved remarkable achievements in the field of depth estimation of panoramic images.
-However, current panoramic depth estimation methods excessively prioritize local pixel accuracy over geometric room structures, leading to two issues: (1) Noisy ground truth data induces local overfitting that propagates errors to adjacent regions; (2) Local correlations cause inaccurate predictions at 3D discontinuity regions (e.g., wall corners).
-As described in Fig.
-1, the pre-trained panorama depth estimator PanoFormer[10] shows limited performance on the wall corners while easily be disturbed by noisy ground truth data.
-These two problems caused by ignoring room geometry seriously hinder the application of existing depth estimation models to real-world indoor scenarios.
-To address existing problems in panoramic depth estimation, this paper proposes a depth estimation framework named as room geometry constrained depth estimation network (RGCNet).
-The proposed RGCNet based on multi-task learning unifying three complementary tasks: depth estimation, room layout estimation, and background segmentation.
-This architecture employs a shared panoramic encoder to extract multi-scale features, with three dedicated decoders generating: coarse depth map, room layout map, and background segmentation mask from these features.
-The depth estimation branch retains the architecture of existing high-performance depth estimators [10, 11].
-However, our RGCNet treats the output of the depth decoder as a first-stage prediction and refines it using the room geometry constrain which generated from room layout and background segmentation.
-In our framework, the initial depth estimate and room layout are jointly utilized to compute the background depth maps.
-By incorporating geometric constraints of the room, these derived background depth maps are not only more robust to noise in the ground-truth depth but also yield higher accuracy in structured regions such as walls and corners.
-Furthermore, RGCNet leverages an end-to-end multi-task framework to decode the background depth and fuse it with the coarse depth prediction in a single efficient step.
-To generate the background map, a proposed room-geometry-based background depth resolving strategy leverages the layout and depth decoder’s output to calculating the corresponding background depth.
-To fuse the background and depth decoder’s prediction, a background-segmentation-guided fusion strategy derives fusion weights for the background and coarse depth maps from the segmentation decoder’s predictions.
+Figure 1: This figure presents 3D visualizations of the ground truth depth maps for two rooms with an atypical structure, shown from three distinct viewpoints: frontal, side, and bird’s-eye view (BEV).
+The first row depicts a room whose overall volume approximates a triangular prism shape.
+The second row illustrates a KTV room, in which the sofa cushions and the entire wall are merged into a single continuous structure.
+Existing methods for monocular panoramic depth estimation primarily adopt three technical strategies to address inherent object distortions[10].
+One line of work[25, 38] designs specialized feature extractors, informed by panoramic imaging principles, to directly process distorted equirectangular images.
+Another common strategy [32] projects the equirectangular panorama (ERP) onto another projection, extracts features separately from both, and then fuses them to predict the depth.
+A more recent direction[6] explicitly incorporates geometric priors; it uses estimated room layout to construct a background depth model, which significantly improves accuracy in geometrically discontinuous regions.
+While these methods achieve strong performance on benchmarks with regular, Manhattan-aligned layouts, they often overlook the prevalence of complex shaped rooms in real-world scenarios.
+As illustrated in Fig.
+1, real-world rooms are often characterized by irregular, non-Manhattan layouts.
+In contrast, the background depth estimated by BGDNet [6] based on its predicted room layout typically assumes a regular room structure and known camera pose, which limits the method’s performance in the scenarios depicted in Fig.
+1.
+Nevertheless, prior work [31] has suggested that any irregular room can be decomposed into regular enclosed regions and irregular regions.
+The room layout annotations provided by various datasets usually correspond to the regular enclosed regions in such scenarios.
+Hence, the reliable identification and separation of regular and irregular regions in specifically structured indoor scenes, as well as the subsequent reconstruction of background depth for regular enclosed regions without external measurements, remains an open challenge.
+To address existing problems in panoramic depth estimation, this paper proposes a depth estimation framework named as pose-aware and room geometry-constrained depth estimation framework (PAGCNet).
+The proposed framework based on multi-task learning unifying four tasks: depth estimation, camera pose estimation, room layout estimation, and region segmentation.
+Specifically, region segmentation decoder perform two semantic segmentation task to extract the irregular region mask and background mask from input panorama.
+The irregular region mask used as an indicator to tell depth estimator which region fall outside the enclosed background region defined by room layout, while background mask used to tell our framework to what extent the geometry-constrained background depth should correct the depth decoder’s prediction.
+Building on this output, our PA-BDR component optimize the camera pose decoder’s prediction and further resolving the enclosed region background depth.
+Then, our FGM component produces the fusion weight map according to the two mask prediction of region segmentation decoder.
+Finally, an adaptive fusion component integrates the enclosed region background depth with the depth decoder’s prediction under the guidance of fusion weight map, yielding a refined depth prediction.
+Extensive experimental results on the Matterport3D, Structured3D, Replica dataset show that our proposed methods can achieve significantly superior performance than current open-source methods.
 In general, our contributions can be summarized into the following three points:
-  • A room-layout constrained depth estimation framework RGCNet has been proposed to use room structural geometry to get corrected depth predictions.
-  • A background depth map resolved strategy has been proposed, which extract the room structural geometry information from multi-task decoder’s prediction.
-  • A adaptively fusion strategy has been proposed, which adaptively based on extracted the geometry information to refine the depth prediction.
+  • We proposes PAGCNet, which calculates the enclosed region background depth to adaptively optimizes the final depth prediction.
+  • We design a PA-BDR component to compute the background depth without external measurement of camera pose.
+  • We introduce an FMG component and an adaptive fusion component.
+The FMG component uses the region segmentation decoder’s output to compute the fusion weight, while the adaptive fusion component uses this fusion weight to perform the fusion between background depth and depth decoder’s prediction.
 II Related Works
@@
 Processing typically involves projecting these spherical images onto 2D planes through perspective mapping.
-Common modalities include equirectangular[13, 14, 10, 11, 15, 16, 17], tangent[18, 19, 20], and icosahedron projections[21, 22].
+Common modalities include equirectangular[44, 7, 25, 38, 40], tangent[8, 18, 24], and icosahedron projections[39, 15].
 For depth estimation, equirectangular projection predominates.
 Omnidepth[44] introduced RectNet for efficient equirectangular feature extraction and depth prediction.
-ODE-CNN[14] proposed a hardware-software co-design system comprising: 1) a panoramic camera with binocular depth sensors, and 2) a depth estimation model leveraging RectNet baseline enhanced with spherical feature transformers between encoder-decoder stages and deformable convolutional spatial propagation for final prediction.
+ODE-CNN[7] proposed a hardware-software co-design system.
 Following vision transformers’ emergence, specialized models have proliferated: Panoformer[25] developed reference-point window self-attention for equirectangular features within an encoder-decoder architecture; Egformer[38], SGFormer[40], GLPanoDepth[3] subsequently incorporated global receptive fields and spherical geometry constraints.
-To address panoramic depth data scarcity, self-supervised methods [16, 17] operate directly on equirectangular RGB images.
 Alternatively, tangent-view approaches [8, 18, 24] project panoramas onto multiple views, estimate per-view depth, and spatially composite patches into panoramic depth maps.
@@
 To address this limitation, recent approaches[6] integrate room structural priors into panoramic depth estimation.
-The core challenge lies in efficiently computing accurate background depth maps, which require precise room layout estimation.
-Indoor panoramic layout estimation aims to detect room wall boundaries from input panoramas.
-Existing research predominantly adopts the Manhattan World assumption.
-LayoutNet[33] directly predicts per-pixel corner and boundary probability maps.
-Conversely, Dula-Net[34] decodes panoramas into equirectangular and perspective views, extracting wall-floor/ceiling boundaries via semantic segmentation.
-EquiConv[35] employs specialized convolutions to generate corner/edge probability maps.
-These early methods share a 2D segmentation paradigm, outputting full-resolution probability maps—an inefficient approach given layout estimation only requires boundary/corner identification.
-HorizonNet[36] thus reformulates the task as 1D sequence prediction using Bi-LSTMs, while HoHoNet[29] employs multi-head self-attention for similar 1D representation.
-LED-Net[37] further predicts layouts from horizontally compressed representations.
-Subsequently, AtlantaNet[38] projects panoramas onto ceiling/floor planes to avoid occlusion, predicting layouts via amalgamated projections.
-DMH-Net[39] extends this by mapping to cubemap faces before predicting boundaries and corners.
-A critical secondary challenge involves fusing background depth with panoramic depth predictions.
-The recent BGDNet[32] incorporates room geometry as background depth when predicting the final depth map.
-However, it needs pretrained HorizonNet and SAM[40] to extract the background depth map from input panorama image.
-In this paper, we adopt a multi-task learning approach to integrate room layout and segmentation predictions needed for background depth calculation within a single framework.
-We also design a background depth calculation strategy based on background segmentation and a fusion strategy for coarse and background depth guided by background segmentation weights.
+Driven by advances in room layout estimation, current models[28, 42, 23, 33, 1, 31] demonstrate high accuracy in predicting layouts for rooms with regular, Manhattan-aligned structures.
+Yet, in real-world indoor images, room structures are often irregular and not based on a Manhattan layout, and the camera’s pose itself cannot be obtained in real time.
+This severely limits its performance in real-world indoor scenes with irregular structures.
+To eliminate the dependence on these two assumptions, our multi-task learning framework incorporates camera pose learning and irregular region segmentation tasks.
 Our method is detailed in the following sections.
-Figure 2: The structure diagram of our proposed room geometry guided depth estimation framework.
-In terms of model structure, our framework includes a shared panorama encoder and three task-corresponding decoders.
-Based on the obtained layout map, coarse depth map, and background-segmentation map, the proposed framework decode fine-grained depth prediction.
+Figure 2: The structure diagram of our proposed PAGCNet.
+In terms of model structure, our framework includes a shared panorama encoder and four task-corresponding decoders.
+Based on the obtained layout estimation, first stage depth, camera pose, irregular region mask and background region mask, the proposed framework decode fine-grained depth prediction.
 III Methodology
-The room structural regularized depth estimation framework proposed in this paper is shown in Fig.2.
-From the view of neural network architecture, the framework includes a shared panorama encoder and three task-corresponding decoder modules.
-Moreover, our framework contain a room geometry-based background depth resolving strategy and a background-segmentation-guided fusion mechanism.
+The proposed PAGCNet’s diagram in this paper is shown in Fig.2.
+From the view of network architecture, the framework includes a shared panorama encoder and four task-corresponding decoder modules.
+Moreover, our framework contain a PA-BDR component, a FGM component and an adaptive fusion component.
 The panorama encoder extract the multi-stage feature sets $F=\{f^{2^{i+1}\times}_{i}|i=1,2,3,4\}$ from the panorama image.
-The decoder modules corresponding to the subsequent three tasks use these features as input to estimate the corresponding results.
-The proposed room geometry-based background depth resolving strategy calculate the background depth based on panorama depth, room layout, background segmentation predictions.
-Then, background-segmentation-guided fusion mechanism regularize the depth decoder’s prediction based on background depth map and background segmentation predictions.
+The proposed PA-BDR component optimize the camera pose and resolving regular enclosed region’s background depth based on tasks decoder’s prediction.
+Then, FMG component extract the fusion weight map from input panorama with two sub binary semantic segmentation task’s prediction.
+Finally, the adaptive fusion component will fuse the regular enclosed region’s background depth and coarse depth prediction of depth decoder to get the refined depth predictions.
 In the following subsections we will introduce each module in detail.
 III-A Architecture
-Figure 3: The structure diagram of layout feature aggregation module and pixel feature aggregation module.
-Panorama Encoder: In this paper, the framework we proposed uses the backbone proposed by PanoFormer as a common feature encoder.
+Panorama Encoder: In this paper, the framework we proposed uses the backbone proposed by PanoFormer[25] as a common feature encoder.
 The backbone consists of the panorama transformer block and convolution 2D layer based downsample layer.
-The panorama transformer block contains a window self-attention mechanism designed for the panorama imaging process and a feed-forward layer designed based on depthwise separable convolution[41].
+The panorama transformer block contains a window self-attention mechanism designed for the panorama imaging process and a feed-forward layer designed based on depth-wise separable convolution[12].
 The entire backbone consists of an input projection layer and $4$ stage blocks.
-The input projection layer consists of a convolution layer whose kernel size is $3\times 3$ with stride of $2$, a 2D batchnorm layer and a ReLU activation function.
+The input projection layer consists of a convolution layer whose kernel size is $3\times 3$ with stride of $2$, a 2D batch-norm layer and a ReLU activation function.
 The subsequent $4$ stage blocks all consist of a panorama transformer block and a double-downsampling layer composed of a 2D convolutional network.
 Each stage block will output the corresponding double-downsampled feature map result.
-Layout Decoder: Considering that layout estimation is a non-pixel-level task while depth estimation and background segmentation are both pixel-level tasks, and some existing work [42] also mentioned that different tasks may have different requirements for features of different scales.
-To this end, we design a specific layout FA module for the layout task to further process the features extracted by the panorama encoder.
-The structure diagram of the layout feature aggregation module for the layout task is shown in Fig.3 (a) .
-This module use the height compress convolution with spatial stride $(2,1)$ to compress the feature maps $f^{4\times}_{1},f^{8\times}_{2},f^{16\times}_{3},f^{32\times}_{4}$.
-Then the compressed feature map is flattened to obtain a feature vectors set $\hat{f}_{1},\hat{f}_{2},\hat{f}_{3},\hat{f}_{4}$.
-We then cat these flattened features together and, following the Hohonet[25] model design experience, use a multi-head self-attention module to extract global feature information from the flattened feature vectors and output the final feature vector $\hat{f}_{cat}$.
-Based on the output features of the layout FA module, the layout decoder uses boundary and corner heads, consisting of a convolution 1D layer, a ReLU layer, and a batchnorm 1D layer, to predict the final room layout $S_{room}$.
-Depth Decoder: Considering that both depth estimation and background segmentation tasks are pixel-level tasks, we follow the design idea of Hohonet and let them use a common pixel-level feature aggregation to process the multi-stage panorama feature.
-The structure of the pixel-level feature aggregation module we use is shown in Fig.
-3 (b).
-This module consists of a panorama transformer block and an upsampling module.
-Given multi-scale feature map set $f^{1}_{4\times},f^{2}_{8\times},f^{3}_{16\times},f^{4}_{32\times}$, this module’s feature integration process can be described as:
-$\hat{f}=\prod^{3}_{i=1}[UP(PT(f^{2^{i}\times}_{i}))+f^{2^{i+1}\times}_{i+1}]$ (1)
-On the integrated multi-scale feature $\hat{f}$, the depth estimation decoder predict the coarse depth maps $S^{p}_{depth}=\{d_{ij}|i=1,2,...,W;j=1,2,...,H\}$.
-background-Segmentation Decoder: As mentioned above, the background segmentation task and the depth estimation task share a pixel-level feature aggregation module.
-The structure diagram of the pixel-level feature aggregation module is shown in Fig.
-3(b).
-Therefore, in our framework, the background-segmentation decoder itself has only one prediction head consisting of a convolution 2D layer, a ReLU function, and a batchnorm 2D layer.
-This prediction head predicts the corresponding 2D background segmentation result $S_{seg}=\{0\leq p_{ij}\leq 1|i=1,2,...,W;j=1,2,...,H\}$ based on the integrated multi-scale feature $\hat{f}$.
-III-B Room Geometry-based Background Depth Resolving Strategy
+Layout Decoder: Considering that layout and camera pose estimation are non-pixel-level task while depth estimation and region segmentation are both pixel-level tasks.
+Some works[41] mentioned that different tasks may have different requirements for features of different scales.
+Following the hohonet[29], we first employ a height compression module to project multi-scale panoramic features from 2D into a 1D sequential representation.
+This sequence is then processed by the Transformer encoder introduced in Hohonet[29], and finally passed through a lightweight MLP to predict the enclosed room layout $S_{room}$.
+Camera Pose Decoder: The camera pose decoder follows a feature processing pipeline consistent with that of the layout decoder: multi-scale 2D panoramic features are first compressed along the height dimension into a 1D sequence, which is then processed by a Transformer encoder.
+However, due to the distinct nature of the camera pose estimation task, the corresponding feature processing blocks in this decoder do not share weights with those in the layout decoder.
+After feature encoding, the camera pose is regressed through a lightweight multi-layer perceptron (MLP) head.
+Depth Decoder: In this section, we maintain consistency with the original PanoFormer[25] paper.
+This part of the network consists of panorama transformer blocks and linear interpolation upsampling layers.
+The decoder network is structurally symmetrical to the panorama encoder.
+On the last panorama transformer block’s feature output $\hat{f}$, the depth estimation decoder predict the coarse depth maps $S^{p}_{depth}=\{d_{ij}|i=1,2,...,W;j=1,2,...,H\}$.
+Region Segmentation Decoder: Given that the panorama encoder and depth decoder together constitute a U-Net-like architecture—commonly employed in semantic segmentation—we do not introduce separate feature extraction modules for the region segmentation decoder.
+Instead, we directly utilize the feature map $\hat{f}$ output by the final transformer block in the depth decoder.
+Two independent convolutional layers are then applied to this feature map to predict the irregular region mask $S_{ir}=\{p_{ij}\in{0,1}|i=1,2,...,W;j=1,2,...,H\}$ and the background segmentation mask $S_{seg}=\{0\leq p_{ij}\leq 1|i=1,2,...,W;j=1,2,...,H\}$, respectively.
+III-B Pose Aware Background Depth Resolving Component
 Figure 3: P is the camera center, A and B are the upper and lower boundary points of the wall corresponding to point P, and D is an arbitrary point on the wall plane.
@@
 Based on the depth of point P predicted by the depth decoder, $d_{c}$ and $d_{f}$ can be calculated.
-Given a room layout estimation prediction $S_{room}\in R^{W\times 3}$ to resolve the corresponding background depth, we need to obtain the camera-to-ground distance information.
-Existing methods often assume a fixed camera height when calculating background depth based on layout information.
+III-B1 Camera Height Optimize
+Given a room layout estimation prediction $S_{room}\in R^{W\times 3}$ to resolve the corresponding background depth, we need to obtain the camera-to-ground and camera-to-ceiling distance information.
+Existing methods often assume a fixed camera height or camera height can be obtained easily.
 This assumption is stable and reliable in virtual rendering datasets, but it is not necessarily reliable when collecting data in the real world using a panoramic depth camera.
-To obtain more stable and reliable camera height information for each scene, we first designed a camera height resolution strategy.
-As shown in the Fig.
-4, define the coordinates of the camera origin $O$, and draw a normal line from the camera plane to the image plane, which intersects the wall plane at point $P$.
-Then draw a straight line perpendicular to the camera plane at point $P$, which intersects with the ceiling-wall boundary line in the room layout prediction at point $A$, and intersects with the floor-wall boundary line at point $B$.
-What our camera height calculation strategy needs to do is to calculate the camera height in the Fig.
-4.
-Our method use the predictions of the depth decoder and room layout decoder as the input of camera height calculating method.
-The room layout predictions are responsible for providing the image plane coordinates of the intersection points $A$ and $B$ in the panorama image and depth maps.
-Based on the results of the depth maps prediction, we can get the distance between the corresponding line segments $OA$ and $OB$.
-On this basis, the distance from the camera to the ceiling and the camera height can be conveniently calculated using the following formula:
-$|AP|=d_{c}\times sin(\phi_{c})$ (2)
-$|PB|=d_{f}\times sin{\phi_{f}}$
-The corresponding angles $\phi_{c}$ and $\phi_{f}$ can be calculated using the geometric imaging principle of a panoramic camera.
+To obtain a stable and reliable camera height, our framework first use a camera pose decoder to predict an init camera height $\hat{h}_{c}$ from input panorama.
+Then, the PA-DBA component use the depth, room layout decoder’s prediction to calculate a camera height $\tilde{h}_{c}$.
+Finally, we count the mean value of $\hat{h}_{c}$ and $\tilde{h}_{c}$ to solve the accurate camera height $h_{c}$.
+In the following, we show the detail of calculation process of $\tilde{h}_{c}$.
+From the results of room layout and region segmentation, we can obtain the foreground pixel coordinates of the upper and lower boundaries of the wall without occlusion.
+Here, we take a straightforward and easy-to-understand example (as the Fig.
+3) to explain how to calculate our camera height $\tilde{h}_{c}$ from these occlusion free wall boundaries pixel coordinates.
+As shown in Fig.
+3, we define the origin of the camera plane as $O$, and then draw a horizon line on the camera plane centered at point $O$.
+Based on the results of the region segmentation decoder, we can select the portion of horizon line without occlusion and falling in the wall.
+We randomly select a point $P$ from these segment regions; the length of $OP$ is then the depth value of the corresponding pixel in the coarse depth map.
+Simultaneously, based on the horizontal coordinates of point $P$, we can obtain the two points $A$ and $B$ above and below the corresponding wall boundary line.
+The lengths of $OA$ and $OB$ can also be obtained from the corresponding depth map pixels $d_{c}$ and $d_{f}$.
+Using basic geometry, we can calculate the camera height $AP$ and the height $BP$ from the camera to the ceiling.
+The calculation formulas are as follows:
+$|AP|=d_{c}\times sin(\phi_{c}),$ $\ |PB|=d_{f}\times sin{\phi_{f}},$ (1)
+where angles $\phi_{c}$ and $\phi_{f}$ can be calculated using the geometric imaging principle of a panoramic camera.
 Assuming that the vertical coordinate of point $A$ is $u_{ceil}$ and the vertical coordinate of point B is $u_{floor}$, the calculation formulas for the two angles can be obtained using the following formulas:
@@
 […]
-After determining the camera height and the distance between the camera and the ceiling, we can use the wall information from the room layout predictions to calculate the corresponding background depth maps for each pixel belonging to the background wall.
+Through the above process, we can obtain a set of camera heights and distances from the camera to the ceiling.
+However, this calculation process is greatly affected by depth estimation errors.
+To reduce this impact, we use the horizontal line passing through point $O$ as a reference and perform the above calculation process at each occlusion-free point $P$ on this line segment.
+Finally, we average all the calculated camera heights to obtain the final $\tilde{h}_{c}$.
+Ultimately, averaging $\hat{h}_{c}$ and $\tilde{h}_{c}$ yields a relatively accurate camera height $h_{c}$.
+III-B2 Enclosed Background Depth Resolving
+After determining the camera height $h_{c}$ and the distance between the camera and the ceiling, we can use the wall information from the room layout predictions to calculate the corresponding background depth maps for each pixel belonging to the background wall.
 The depth maps for the ceiling and floor can be simply calculated based on the angles calculated from the pixels using the following formula:
-$d^{i}_{floor}=\frac{|PB|\times H}{(u^{i}_{floor}-0.5H)\times\pi}$ (4)
-$d^{i}_{ceil}=\frac{|AP|\times H}{(0.5H-u^{i}_{ceil})\times\pi}$
+$d^{i}_{floor}=\frac{h_{c}\times H}{(u^{i}_{floor}-0.5H)\times\pi},$ (3)
+$d^{i}_{ceil}=\frac{|AP|\times H}{(0.5H-u^{i}_{ceil})\times\pi},$
 For any point $D$ on the wall plane that is tilted toward the floor, assume its equirectangular plane coordinates on the panorama image are $(u_{i},v_{i})$.
@@
 […]
-$\rho_{D}=(1-\frac{v_{i}}{W})\times\pi$
+$\rho_{D}=(1-\frac{v_{i}}{W})\times\pi,$
 Based on the two deflection angles, to obtain the depth of $D$, $d_{D}=|OD|$, we first draw a perpendicular line through point $D$ to line segment $AB$ and intersect $AB$ at point $E$.
@@
 Finally, combining all the solved depth values, we can obtain $S_{back}$.
-III-C Background-segmentation-guided Fusion Mechanism
-As mentioned above, based on the corase depth map and the background depth map, our framework needs to fuse the predicted depth $S^{p}_{depth}$ and the background depth $S_{back}$ to obtain an accurate final depth estimation result $S^{final}_{depth}$.
-Considering that $S_{back}$ is naturally suitable as an upper bound for depth estimation when the layout prediction is relatively accurate.
-Then $S_{back}$ as an upper bound can be used to ensure that the depth of pixels originally belonging to the rescue, ceiling, ground and other areas do not exceed the area of the room itself.
-The depth value of the pixel belonging to the foreground object area can be as close to $S^{p}_{depth}$ as possible.
-Therefore, in the process of fusing the two depth maps, we need a weight to determine whether the current pixel is an object belonging to the foreground area or the background area.
-The probability map $S_{seg}$ predicted by the background-segmentation task in our framework can just be used as such a fusion weight.
-Therefore, the entire fusion process can be described as the following formula:
-$S^{final}_{depth}$ $=\{d^{final}_{ij}|i=1,2,.,W;j=1,2,.,H\};$ (6)
-$etl.$ $\ d^{final}_{ij}=d^{back}_{ij}\times p_{ij}+d^{p}_{ij}\times(1-p_{ij});$
-$p_{ij}\in S_{seg};$
-$d^{back}_{ij}\in S_{back};$
-$d^{p}_{ij}\in S^{p}_{depth};$
-III-D Objective Function
-[…]
-When training the layout branch, we refer to the settings of HorizonNet and use the binary cross entropy and L1 loss function to calculate.
-When training the depth estimation branch, we used the same Huber[43] (or Berhu[44]) loss and gradient loss[45] as the baseline Panoformer.
-The hyper-parameters within the objective functions of these two branches were also consistent with those of HorizonNet and PanoFormer.
-Considering that existing panoramic image datasets usually do not contain background segmentation annotations.
-When training the background segmentation decoder, we first use the background depth map $S_{back}$ predicted by the framework and the corresponding ground-truth depth map $S^{gt}_{depth}$ to calculate the corresponding background segmentation label $S^{gt}_{seg}$.
-The specific solution formula is as follows:
-$D^{res}=|S^{gt}_{depth}-S_{back}|;$ (7)
-$S^{gt}_{seg}=\{p^{gt}_{ij}=\Gamma(d^{res}_{ij}<\gamma)|d^{res}_{ij}\in D^{res}\}$
-where $\Gamma$ is a binary function, which takes the value 1 when the condition in the brackets is true and takes the value 0 when it is false.
-When optimizing the partial network for the background-segmentation task, we use a simple focal loss as our object function, which is calculated as follows:
-$L_{seg}=-\frac{1}{M}\sum^{M}_{u=1}\frac{1}{H*W}\sum^{W}_{i=0}\sum^{H}_{j=0}\alpha(1-\hat{p}_{ij})^{\eta}log(\hat{p}_{ij});$ (8)
-where $\hat{p}_{ij}=p_{ij}$ if $p^{gt}_{ij}=1$ else $\hat{p}_{ij}=1-p_{ij}$. Following the common setting of focal loss, we set the $\alpha$, $\beta$ as 0.5, 2.0 relatively.
-In summary, the overall object function of our proposed framework can be described as:
-$L_{all}=\lambda_{1}*L_{layout}+\lambda_{2}*L_{depth}+\lambda_{3}*L_{seg}$ (9)
-where $\lambda_{1},\lambda_{2},\lambda_{3}$ are three manual setting parameters.
+III-C Fusion Mask Generation Component
+A crucial step in constraining the depth decoder’s prediction using the calculated background depth is identifying which regions in the input panoramic image are valid.
+Previous work, BGDNet, simply used the semantic segmentation mask output by the SAM model as the criterion for this.
+This approach effectively distinguishes which parts of a room’s interior are foreground objects and which areas belong to the background walls.
+However, this approach relies on two prerequisites: 1) no large areas in the input image extend beyond the room structure; and 2) the room structure itself is regular and conforms to the Manhattan layout.
+Previous work has mentioned that any irrationally shaped room structure can essentially be divided into regular and irregular regions.
+The room layouts typically provided by datasets describe enclosed regular regions within the room structure.
+Therefore, to make the background depth-based method more generalizable, we need to determine not only whether a pixel belongs to the foreground object but also whether the region belongs to the regularly enclosed region defined by the room layout.
+In our framework, we treat the distinction between foreground and background regions, and between regular enclosed regions and irregular regions, as two independent binary semantic segmentation tasks.
+Both of these semantic segmentation tasks are performed by our region segmentation decoder.
+This decoder outputting the irregular region mask $S_{ir}$ and the background mask $S_{seg}$.
+In $S_{ir}$, each element $s^{ir}_{uv}$ represents the probability that the pixel at row $u$ and column $v$ in the input panorama belongs to an irregular region.
+Similarly, in $S^{seg}$, each element $s^{seg}_{uv}$ denotes the probability that the pixel belongs to the background.
+To properly fuse coarse depth prediction and background depth, we need to calculate a fusion weight to balance their respective proportions in the final output.
+Here, we first binarize each element in the region mask $S_{ir}$ using a threshold of $0.5$.
+Then, we use the sigmoid function to compress the value of each element in $S_{seg}$ to the range $[0,1]$.
+Finally, we multiply each element in $S_{ir}$ by each element in $S_{seg}$ to obtain the final fusion weight mask $S_{weight}$.
+III-D Adaptive Fusion Component
+As mentioned above, based on the depth decoder’s prediction and the enclosed background depth map, our framework needs to fuse the predicted depth $S^{p}_{depth}$ and the enclosed background depth $S_{back}$ to obtain an accurate final depth estimation result $S^{final}_{depth}$.
+Considering that $S_{back}$ is naturally suitable as an upper bound for depth estimation when the layout prediction is relatively accurate.Then $S_{back}$ as an upper bound can be used to ensure that the depth of pixels originally belonging to the rescue, ceiling, ground and other areas belonging to regular regions do not exceed the area of the room itself.
+And, the depth value of the pixel belonging to the foreground object or irregular region can be as close to $S^{p}_{depth}$ as possible.
+The entire fusion can be described as following formula:
+$S^{final}_{depth}$ $=\{d^{f}_{uv}|u=1,2,.,W;v=1,2,.,H\};$ (5)
+$etl.$ $\ d^{f}_{uv}=s^{uv}_{weight}\times d^{b}_{uv}+(1-s^{uv}_{weight})\times d^{p}_{uv},$
+$s^{uv}_{weight}\in S_{weight};d^{b}_{uv}\in S_{back};d^{p}_{uv}\in S^{p}_{depth},$
+III-E Objective Function
+Our framework as a whole includes four task decoders: layout estimation, depth estimation, camera pose estimation, region segmentation.
+The objective functions corresponding to the four decoders can be expressed as $L_{layout}$, $L^{p}_{depth}$, $L_{region}$, and $L_{pose}$, respectively.
+Since most existing datasets lack comprehensive annotations for room layout estimation, we first aggregate all layout-related annotations from the Structured3D and Matterport3D datasets to form a large-scale layout dataset.
+Following the practice described in[31], we apply a simple script to convert the existing room layout annotations—which may contain occlusions or ambiguities—into regular and enclosed layouts.
+Using this generated dataset, we first pre-train the room layout branch of our proposed framework.
+For this stage, similar to HorizonNet, we adopt a combination of binary cross-entropy loss and L1 loss.
+For the depth estimation branch, we follow Panoformer and employ a Huber[9] loss together with a gradient loss[26], keeping the hyper-parameter settings consistent with those in the original HorizonNet[28] and Panoformer papers respectively.
+For the two semantic segmentation subtasks, we use the commonly adopted binary cross-entropy loss and Dice loss[17].
+As for the camera pose estimation, we simply apply a smooth L1 loss.
+Overall, the total loss function of our proposed framework can be expressed as:
+$L_{all}=\lambda_{1}*L_{layout}+\lambda_{2}*L_{depth}+\lambda_{3}*L_{region}+\lambda_{4}*L_{pose},$ (6)
+where $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ are four manual setting parameters.
 IV Experiment
@@
 In this section, we introduce the experiments to verify the proposed depth estimation framework.
-In our experiments, we selected three datasets: Stanford2D-3D[46], MatterPort3D[47], and Structure3D[48].
-Stanford2D-3D and MatterPort3D are two datasets collected in the real world, while the Structure3D dataset is a synthesized dataset.
-Real-world datasets contain a large amount of noise points introduced during the acquisition process as described in OnimiDepth[13].
-This noise introduces local jumps that severely hinder model performance.
-Therefore, before training models on Stanford2D3D and Matterport3D, we used a proposed layout-based dataset denoise strategy to constrain the depth of these noise points to within the room.
-Stanford2D-3D: The Stanford2D-3D dataset contains 1413 panoramic images collected from 3 types of buildings divided into 6 large areas.
-We follow the official practice to divide the dataset into training and test sets, and downsample all depth maps and RGB images to $512\times 1024$ size images.
+In our experiments, we selected three datasets: MatterPort3D[5], Structure3D[43], Replica[30].
+MatterPort3D and Reiplica is collected in the real world, while the Structure3D dataset is a synthesized dataset.
 Matterport3D: Matterport3D contains 10,800 panoramic images collected from 90 different rooms.
@@
 In this dataset, we follow the official recommended setting [43], using the data of the first 3000 scenes as the training set, the data of 3000-3250 scenes as the validation set, and the data of the last 250 scenes as the test set.
-Metric: Following the previous works, we use some standard evaluation metrics, which include: relative error (abs rel), squared relative error (sq rel), root mean squared error (rmse), and three threshold percentage $\delta<\varsigma^{t}(\varsigma=1.25,t=1,2,3)$ denoted as $\delta_{t}$.
-Dataset denoise strategy: Considering that datasets collected in the real world contain a large amount of noisy points, we have also designed a denoising strategy for panoramic depth map datasets.
-As mentioned earlier, our method uses the outputs of the layout decoder and depth decoder to calculate the corresponding background depth map.
-Here, we use this calculated background depth map to constrain the dataset collected in the real world.
-First, for areas in the real-world depth ground-truth maps where measurement fails, we directly replace them with the background depth value.
-Then, our method sets a threshold to judge the ground-truth depth.
-If the depth point in the ground-truth dataset of the original data is converted to 3D space and its coordinates are outside the room and the distance from the room wall is more than 1 meter, we identify the pixel as a noise pixel.
-For these pixels, we also use the depth in the background depth map as a replacement.
+Replica: The Replica dataset is captured using an RGB-D camera rig with an IR projector from real-world indoor scenes.
+We follow the BGDNet[6] render 3554 images from these 3D models by setting camera in random position on the floor.
+Metric: Following the previous works[3, 25, 40], we use some standard evaluation metrics to measure our methods, which include: mean relative error (MRE), mean average error (MAE), root mean squared error (RMSE), and three threshold percentage $\delta<\varsigma^{t}(\varsigma=1.25,t=1,2,3)$ denoted as $\delta^{t}$.
+For the room layout task, we used 2D IoU, 3D IoU, RMSE, and $\delta^{1}$ for evaluation.
+For the camera pose task, we used relative rotation error(RRA), relative translation error(RTA), and area under the curve(AUC) for evaluation, with a threshold of 30 for each metric.
+For the region segmentation task, we used mIoU for evaluation.
 Training setting: We use Adam as the optimizer, and the parameters of the optimizer are basically the basic settings of the pytorch framework.
@@
 For horizontal angle rotation, we set the interval of random angle to $\frac{\pi}{4},\frac{\pi}{4}]$.
-Parameter setting: The training objective function used by our framework is shown in Section 4.4, which contains three subcomponents: $L_{front}$, $L_{depth}$, and $L_{layout}$.
-Considering that some datasets do not contain complete layout annotations, and our framework is extremely dependent on the accuracy of layout estimation, we will first extract all layout annotations from the three selected datasets to form a large layout dataset to pretrain the layout task.
-Then, when training the depth estimation task on specific datasets, we decide whether to finetune the weights of the layout and feature encoder parts based on whether the dataset contains complete layout annotations.
-Specifically, the Stanford2D-3D data does not provide complete layout annotations, so we set the corresponding hyper-parameter $\lambda_{1}$ to 0 when training our framework, and the corresponding $\lambda_{2},\lambda_{3}$ to 1.0 and 0.4.
-The Matterport3D and Structure3D datasets provide complete layout annotations, so we set $\lambda_{1},\lambda_{2},\lambda_{3}$ to 0.01, 1.0, and 0.4, respectively.
-In addition, some hand-tuned hyper-parameter settings related to this study, such as $\rho_{1}$ and $\rho_{2}$ in equation was set to 1.0 and 0.5.
+Parameter setting: The training objective function used by our framework is shown in Section III.
+B , which contains four subcomponents: $L_{pose}$, $L_{region}$, $L_{depth}$, and $L_{layout}$.
+During pre-training of the layout decoder, we set the $\lambda_{1}$ to 0 and other parts to 0, while during the training the training of the other decoders, we set $\lambda_{2},\lambda_{3},\lambda_{4}$ as 1.0, 0.4, 0.6.
 Other hyper-parameters that are not relevant to this study are set according to the common specifications in the current field.
-IV-B Performance Comparison
-Figure 5: The 3D visualize results of final depth estimation.
-For each scene, we selected three perspectives: top view, side view, and internal perspective to display the three-dimensional visualization effect of the point cloud converted from the depth map.
-IV-B1 Quantification Comparison with SoTA Depth Estimation Methods
 TABLE I: Quantification comparison with State of the Art depth depth estimation methods on Matterport3D
@@
 $\delta_{1}$ $\delta_{2}$ $\delta_{3}$ RMSE MRE MAE
-Matterport3D EGFormer[11] ICCV 2023 0.8158 0.9390 0.9735 0.6025 0.1517 0.1473
-OminiFusion[19] CVPR 2022 0.9040 0.9757 0.9919 0.4261 0.0552 0.0900
-Bifuse[12] CVPR 2020 0.8452 0.9319 0.9632 0.6295 0.2408 0.3470
-UniFuse[28] IEEE RAL 2021 0.8897 0.9623 0.9831 0.4941 - 0.2814
+Matterport3D Panoformer[25] ECCV 2022 0.9184 0.9804 0.9916 0.3635 0.0571 0.1013
+EGFormer[38] ICCV 2023 0.8158 0.9390 0.9735 0.6025 0.1517 0.1473
 HRDFuse[1] CVPR 2023 0.9162 0.9669 0.9844 0.4433 0.0936 0.0967
-RSDNet[50] CVPR 2022 0.443 0.097 0.248 0.906 0.971 -
+GLPanoDepth[3] IEEE TIP 2024 0.8641 0.9561 0.9808 0.5223 - 0.2998
+GADFNet[13] IEEE TCSVT 2025 0.9047 0.9668 0.9868 0.4564 0.0963 0.2592
 SN360[21] IEEE Access 2025 0.4483 - - 0.9392 0.9808 0.9932
-GLPanoDepth[23] IEEE TIP 2024 0.8641 0.9561 0.9808 0.5223 - 0.2998
-Panoformer[10] ECCV 2022 0.9184 0.9804 0.9916 0.3635 0.0571 0.1013
 SGFormer[40] IEEE TCSVT 2025 0.8946 0.9642 0.9859 0.4790 - 0.2748
-Ours 0.9199 0.9820 0.9984 0.2436 0.0425 0.0829
-TABLE II: Quantification comparison with State of The Art depth estimation methods on Stanfor2d3d
+DepthAnyPanorama[20] arxiv 2025 0.8518 0.9113 0.9764 0.7510 0.4456 0.7845
+DepthAnyDirection[16] ICLR 2026 0.9561 0.9860 0.9982 0.2882 0.0766 0.0912
+Ours 0.9399 0.9852 0.9984 0.2236 0.0411 0.0629
+TABLE II: Quantification comparison with State of the Art depth estimation methods on Structured3D
 Dataset Method Pub’ Year Classic Metrics
 $\delta_{1}$ $\delta_{2}$ $\delta_{3}$ RMSE MRE MAE
-Stanford2d3d EGFormer[11] ICCV 2023 0.8185 0.9338 0.9736 0.4974 0.1408 0.1528
-OminiFusion[19] CVPR 2022 0.8940 0.9714 0.9900 0.3715 0.0543 0.0961
-Bifuse[12] CVPR 2020 0.8660 0.9580 0.9860 0.4142 0.1209 0.2343
-UniFuse[28] IEEE RAL 2021 0.8711 0.9664 0.9882 0.3691 - 0.2082
-HRDFuse[29] CVPR 2023 0.8941 0.9778 0.9923 0.3452 0.0503 0.0984
-FreDSNet[52] 2022 0.8424 0.9583 0.9863 0.2727 0.0952 0.1327
-RSDNet[50] CVPR 2022 0.394 0.098 0.209 0.903 0.974 -
-SN360[51] IEEE Access 2025 0.2917 - - 0.9369 0.9846 0.9942
-GLPanoDepth[23] IEEE TIP 2024 0.9015 0.9793 0.9901 0.3493 - 0.1932
-Panoformer[10] ECCV 2022 0.9394 0.9838 0.9941 0.3083 0.0405 0.0619
-SGFormer[15] IEEE TCSVT 2025 0.8998 0.9693 0.9908 0.3406 - 0.2017
-Ours 0.9479 0.9857 0.9943 0.2359 0.0285 0.058
-TABLE III: Quantification comparison with State of the Art depth estimation methods on Structured3D
+Structure3d PanoFormer[25] ECCV 2022 0.8943 0.9536 0.97431 0.3017 0.1201 0.1546
+EGFormer[38] ICCV 2023 0.7979 0.9071 0.9455 0.6841 0.4509 0.2205
+HRDFuse[1] CVPR 2023 0.7561 0.9161 0.9631 0.4061 - 0.2451
+SGFormer[40] IEEE TCSVT 2025 0.9613 0.9896 0.9957 0.2429 - -
+GADFNet[13] IEEE TCSVT 2025 0.8647 0.9468 0.9899 0.3524 0.0812 0.1944
+DepthAnyPanorama[20] arxiv 2025 0.8918 0.9243 0.9813 0.6641 0.4456 0.5125
+DepthAnyDirection[16] ICLR 2026 0.9361 0.9713 0.9912 0.2914 0.0686 0.0992
+BGDNet[6] CVPR 2024 0.8336 0.9377 0.9731 0.3490 - 0.1656
+Ours 0.9679 0.9907 0.9983 0.1935 0.0414 0.0613
+TABLE III: Quantification comparison with State of the Art depth estimation methods on Replica
 Dataset Method Pub’ Year Classic Metrics
 $\delta_{1}$ $\delta_{2}$ $\delta_{3}$ RMSE MRE MAE
-Structure3d EGFormer[11] ICCV 2023 0.7979 0.9071 0.9455 0.6841 0.4509 0.2205
-OminiFusion[19] CVPR 2022 0.6921 0.8831 0.9501 0.4951 - 0.2981
-Bifuse[12] CVPR 2020 0.8594 0.9400 0.9672 0.5213 0.2455 0.1573
-UniFuse[28] IEEE RAL 2021 0.8542 0.9399 0.9676 0.5016 0.2319 0.1506
-HRDFuse[29] CVPR 2023 0.7561 0.9161 0.9631 0.4061 - 0.2451
-BGDNet[32] CVPR 2024 0.8336 0.9377 0.9731 0.3490 - 0.1656
-SGFormer[15] IEEE TCSVT 2025 0.9613 0.9896 0.9957 0.2429 - -
-PanoFormer[10] ECCV 2022 0.8943 0.9536 0.97431 0.3017 0.1201 0.1546
-Ours 0.9679 0.9907 0.9983 0.1935 0.0414 0.0613
-Figure 6: Visualization of background depth map guided dataset denoising effect diagram.
-[…]
-The Stanford2d3d dataset is a relatively early dataset, and various existing depth estimation methods have achieved relatively good performance on this dataset.
-However, on this dataset, the noise processing strategy and multi-task depth estimation framework proposed in this paper still show excellent effectiveness.
-Specifically, our method significantly outperforms the current SOTA method by 3.68 percentage points in the RMSE, the evaluation indicator that the industry attaches the most importance to.
-At the same time, in MRE and MAE, which are both numerical indicators, the performance advantage of our method over the existing methods is not as obvious as RMSE, but it is still intuitive, which further proves the robustness and superiority of the method proposed in this paper.
-In the three indicators $\delta_{1},\delta_{2}$ and $\delta_{3}$ at the ratio level, our method also has a leading advantage, but this advantage is not as obvious as the numerical indicators.
-We believe that this is because the existing methods have reached a very high level in the three indicators at the ratio level, so it is more difficult to achieve higher indicators.
+Replica HRDFuse[1] CVPR 2023 0.7711 0.9221 0.9541 0.3891 - 0.1421
+DepthAnyPanorama[20] arxiv 2025 0.9012 0.9335 0.9898 0.4432 0.2121 0.3132
+DepthAnyDirection[16] ICLR 2026 0.9431 0.9668 0.9933 0.2712 0.0733 0.1011
+BGDNet[6] CVPR 2024 0.8554 0.9365 0.9624 0.3456 - 0.1678
+Ours 0.9233 0.9866 0.9992 0.2101 0.0512 0.0629
+IV-B Region Segmentation Decoder’s Annotation Generation
+IV-B1 The generation process of background segmentation
+To efficiently fuse the predictions from the background and depth decoders, we first need information to guide the model in identifying which regions belong to the background and which to the foreground.
+For region segmentation, we refer to BGDNet’s approach, using the corresponding semantic segmentation results to uniformly set all ground, ceiling, and wall surfaces in the scene as the background region.
+Unlike BGDNet, which uses SAM to segment the input image, we directly use the semantic segmentation annotations provided by the dataset to generate the corresponding masks.
+IV-B2 The generation process of irregular region segmentation
+As mentioned in the main text, the background depth generated by BGDNet often ignores the fact that in some scenes, the depth threshold may exceed the background space defined by the room layout, or the entire indoor scene itself may not be a regular Manhattan layout.
+However, non-Manhattan layouts or irregularly shaped rooms will always intrude a regular enclosed Manhattan region, and the room layout annotations provided in the dataset are usually this regular enclosed Manhattan region.
+Therefore, to make this approach of fusing background and depth decoders more generalizable, we need information to help the model distinguish which regions belong to the regular enclosed region.
+Here, to generate a reasonable irregular region segmentation mask, we first calculate the scene background depth for each panel input image using the room layout annotations and camera height from the entire dataset.
+Then, we use the generated background segmentation mask annotation to select the depth of the fused region from both the background depth and ground truth depth annotations.
+Finally, we calculate the difference between the depths of the fused regions selected from the two depths and apply a threshold to obtain the corresponding irregular region segmentation mask annotation.
+In the process of creating the binary mask, we use a threshold of 0.05, which is based on our specific observations in the experiments.
+IV-C Performance Comparison
+IV-C1 Quantification Comparison with SoTA Depth Estimation Methods
+Figure 4: The 3D visualize results of final depth estimation.
+For each scene, we selected three perspectives: top view, side view, front view, and internal perspective to display the three-dimensional visualization effect of the point cloud converted from the depth map.
+In this section, we will conduct performance comparison experiments on three datasets: Matterport3d, Structured3d, and Replica.
+The results have been shown in Table I- III .
+To validate the effectiveness of our approach, we compare our RGCNet with the current state-of-the-art methods[38, 18, 32, 14, 1, 3, 25, 40, 6, 4, 11, 21], including strategies for bi-projection fusion, long-range dependencies, multi-task learning, and background-based methods.
 On the Matterport3D dataset, our method also achieved the current best performance in RMSE, the core evaluation indicator of depth estimation, which fully demonstrated the effectiveness of the proposed method and data processing strategy in key accuracy metrics.
@@
 Despite this, we still achieved a level close to SOTA on this indicator.
-Importantly, this method is significantly ahead of existing work in both the core rmse and rmslog indicators.
+Importantly, this method is significantly ahead of existing work in RMSE.
 On the Structured3d dataset, we mainly compare with BGDNet, which uses a similar idea.
 The results released in the BGDNet paper are trained on the replica and then val on the structure3d dataset, while our method is trained on Structured3D and then validated on the corresponding validation set.
-This comparison may not be fair, but considering that the BGDNet paper is not open source, we can only use this method for comparison.
 Due to the differences in the settings of train and val, the method proposed in this paper has a very obvious advantage in the six indicators of concern.
-TABLE IV: Quatitative analysis of our proposed dataset denoise method.
-Dataset Setting Methods RMSE MRE MAE
-Without Denoise PanoFormer 0.3083 0.0405 0.1013
-Our Methods 0.2673 0.0375 0.0912
-Denoised Dataset PanoFormer 0.2872 0.0392 0.0892
-Our Methods 0.2359 0.0285 0.0580
-TABLE V: The training impact of each task decoder to depth decoder
+On the Replica dataset, we ensured that all data processing steps were as consistent as possible with BGDNet.
+Based on the performance across the six metrics shown, our method significantly outperforms BGDNet.
+Furthermore, it demonstrates a clear advantage over the current state-of-the-art methods, DepthAnyPanorama and DepthAnyDirection.
+It’s worth noting that, considering both of these methods used large datasets to pre-train large models, the performance we present is the result of fine-tuning the officially provided weights on the Replica dataset for two epochs.
+Overall, in terms of quantitative metrics, our method demonstrates a significant performance advantage compared to current state-of-the-art methods, especially in the most common metric rmse.
+This clearly shows that using the background depth we constructed effectively constrains the depth estimator’s predictions, leading to more accurate depth estimation results.
+IV-C2 Visualize Comparison with SoTA Depth Estimation Method
+In this section, we provide a qualitative evaluation of depth estimation results, with visual comparisons shown in Fig.
+4.
+The selected comparison methods include HoHoNet[29], a multi-task learning approach, and EGFormer[38], which is based on a panoramic Transformer architecture.
+The visual results demonstrate that, owing to the effective constraint provided by our background depth modeling, our framework captures the overall geometric structure of the room more accurately than existing open-source state-of-the-art methods.
+This improvement also validates the accuracy of the enclosed-region background depth recovered by our pose-aware depth resolving strategy.
+It should be noted that while our framework predicts the overall room geometry more accurately, its performance remains limited in areas outside the enclosed background region.
+Specifically, in the first scene, distant objects such as the sofa are not well reconstructed.
+In the second scene, certain protruding structures are incorrectly compressed onto the wall planes in the depth prediction.
+The limitation observed in the first scenario is primarily attributed to the fact that the current enclosed-region layout estimation can only represent a limited portion of the overall room structure.
+For depth prediction in areas beyond the enclosed region, our framework relies principally on the capacity of the base depth decoder.
+Nevertheless, accurate depth estimation in such areas remains a persistent challenge for panoramic depth estimators in indoor scenes, and a widely effective solution has yet to be established.
+The issue observed in the second scenario can be attributed to current limitations in irregular region segmentation.
+Inherent inaccuracies in this task lead to certain areas outside the enclosed background being misclassified as background regions.
+Consequently, these areas—which should reflect the room’s extended structure—are incorrectly flattened onto the wall planes in the depth prediction.
+IV-D Quantification Analysis of Background Depth Resolving
+TABLE IV: The performance of our framework’s layout decoder
+Methods 2DIoU 3DIoU RMSE $\delta^{1}$
+HorizonNet[28] 91.18 89.77 0.09 0.99
+HoHoNet[29] 92.96 91.52 0.09 0.99
+Ours 94.36 92.96 0.04 0.99
+TABLE V: The accuracy of our framework’s pose decoder
+Methods RRA@30 RTA@30 AUC@30
+Ours 94.33 90.51 91.34
+TABLE VI: The accuracy of camera height
+mean average error
+camera height from pose decoder $\tilde{h}_{c}$ 0.05
+camera height from depth decoder $\hat{h}_{c}$ 0.04
+camera height from resolving strategy $h_{c}$ 0.02
+TABLE VII: The accuracy of resolved background depth maps.
+Methods RMSE MRE MAE
+HoHoNet’s Background 0.3514 0.1223 0.1923
+PanoFormer’s Background 0.2644 0.0688 0.0991
+Layout GT with Fix Camera pose 0.5514. 0.3846 0.3102
+Layout GT with GT Camera pose 0.0011 0.0010 0.0010
+Layout GT with Pred Camera pose 0.0013 0.0012 0.0012
+Layout Pred with Fix Camera pose 0.5862 0.3212 0.3833
+Layout Pred with GT Camera pose 0.0403 0.0312 0.0213
+Layout Pred with Pred Camera pose 0.0551 0.0417 0.0412
+In this section, we will construct corresponding experiments to verify and compare the performance differences between our background depth calculation strategy and BGDNet’s background depth calculation strategy.
+For convenience, we will only construct experiments on the Structured3D dataset.
+We will first demonstrate the overall performance of the framework’s layout decoder and camera pose parts.
+Based on this, we will show the RMSE, MRE, and MAE of the background region depth and depth ground truth calculated by layout ground truth and layout prediction with and without camera pose.
+Furthermore, we will demonstrate the corresponding metrics of the background depth calculated using our method based on camera pose.
+The results of layout estimation and camera pose estimation in the framework are shown in Tables IV and V, respectively.
+Table VI compares the camera height $\hat{h}_{c}$ calculated from the depth decoder, the camera height $\tilde{h}_{c}$ obtained from the camera pose decoder, and the average of the two $h_{c}$ in our pose-aware background depth resolving strategy.
+The results in Tables IV and V show that the layout decoder and pose estimation results in our framework are quite accurate.
+However, the results in Table 6 show that the camera height calculated from either the camera pose decoder or the depth decoder still has an error of about 5 cm; however, the error of the camera height obtained by averaging the two is reduced to 2 cm.
+This fully demonstrates the effectiveness and rationality of the pose estimation strategy in our proposed method.
+Table VII presents the accuracy of our baseline model in background depth and the experimental results related to background depth calculated based on layout.
+The results show that the depth prediction results of our chosen open-source baseline model are relatively poor in the background portion.
+Furthermore, the background depth value calculated by room layout is highly dependent on the camera height.
+The results in rows 3 and 6 indicate that the camera pose used in generating images in different virtual scenes varies in the structured3d dataset.
+Rows 4 and 7 show that the background depth generated by ground truth layout can be considered, to some extent, as ground truth depth.
+The experimental results in rows 5 and 8 show that the background depth obtained by our proposed background depth calculation strategy also has some minor errors, but these errors are still orders of magnitude different from those produced by other current depth estimators.
+IV-E Quantification and Visualize Analysis of Fusion Mask Generation Component
+TABLE VIII: The mIoU of background segmentation of our region segmentation decoder
+Methods mIoU
+HoHoNet[29] 72.12
+SAN[36] 95.33
+Ours 89.23
+TABLE IX: The mIoU of irregular region segmentation of our region segmentation decoder
+Methods mIoU
+HoHoNet[29] 73.33
+SAN[36] 96.32
+Ours 90.92
+Figure 5: The visualize of resolved background depth.
+The figure shows 3D visualizations of the background depth in four scenes constructed using our method and BGDNet.
+The areas marked with red circles represent regions that BGDNet misjudged as exceeding the background depth, while the areas marked with yellow circles represent regions that our constructed background depth accurately identifies.
+In this section, we construct experiments for the region segmentation decoder and the Fusion Mask Generation Component and verify their performance.
+The dataset used in the experiment is the same as that in the previous subsection, using Structured3D.
+Tables VIII and IX show the performance of the two semantic segmentations performed by the region segmentation decoder.
+In these two subtasks, we chose the open-source HoHoNet as the comparison object.
+We did not choose newer methods because recent methods using multi-task learning on panoramic images are not open-source, so we could not conduct corresponding experiments on our own constructed experimental setup.
+Regarding our Fusion Mask Generation Component, the visualization results of our constructed experiments are shown in Fig.
+5.
+The experimental results in the two tables show that HoHonet exhibits significantly higher performance than the original paper due to the substantial reduction in task difficulty.
+However, considering HoHonet’s overly aggressive approach of compressing 2D features of panoramic images to 1D, its performance still lags considerably behind our method.
+It’s worth noting that while our method represents a significant performance improvement over earlier multi-task learning methods on panorama, it still lags behind recent methods specifically designed for semantic segmentation tasks, such as SAN[36].
+We believe this difference largely stems from our reuse of a portion of the depth decoder network for model overhead considerations.
+As shown in Fig.
+5, BGDNet does not consider that the depth of some regions in panoramic images in real-world scenes may exceed the space defined by the room layout when constructing the background depth.
+This leads to a disastrous problem for BGDNet when constructing the background depth: regions that should extend beyond the background depth are compressed onto the walls defined by the room layout.
+Although the binary semantic segmentation mask provided by SAM used by BGDNet can effectively distinguish which regions are walls, floors, and ceilings, the semantic segmentation mask provided by SAM cannot help the model identify which regions extend beyond the room layout, and these regions also belong to walls, floors, or ceilings.
+In contrast, our method benefits from the irregular region mask provided by the region segmentation decoder, which can effectively distinguish which parts of the scene extend beyond the room layout.
+It is worth noting that since the irregular region mask and background mask themselves are not completely accurate, the background depth constructed by our method has fewer regions for fusion, which reduces the constraint effect of the background depth to some extent.
+However, we believe that reducing some of the blended areas to make the background depth itself more generalizable is a reasonable trade-off.
+IV-F Quantification Analysis of Adaptive Fusion Component
+TABLE X: The comparison of our proposed fusion methods and BGDNet
+Methods RMSE MRE MAE
+Baseline 0.3017 0.1201 0.1546
+Fusion BGDNet[6] 0.2643 0.0812 0.1122
+Fusion ours 0.1953 0.0414 0.0613
+In this section, we construct corresponding experiments to verify the performance of our Adaptive Fusion Component.
+The relevant experimental results are shown in Table X.
+The dataset used in the experiments remains structured3d.
+It is worth noting that, to ensure a fair comparison, we also used the irregular region mask provided by the region segmentation decoder of our framework in the experiments against BGDNet to ensure the accuracy of its background depth.
+As can be seen from the experimental results in Table X, our fusion strategy is significantly better than BGDNet’s approach.
+From the perspective of model structure, the region segmentation decoder in our framework is functionally equivalent to the module in BGDNet that extracts and fuses background depth features, while the subsequent threshold replacement operation in BGDNet functionally corresponds to our adaptive fusion component.
+Based on these analyses, we speculate that the advantage of our fusion method stems from our explicit learning of the weight process for distinguishing foreground and background regions as a supervised learning task during training.
+This learning-based approach is superior to implicitly fusing information into the feature extraction network.
+IV-G Ablation Study
+TABLE XI: The training impact of each task decoder to depth decoder
 Depth
 Decoder
-Front-Seg
+Camera Pose
 Decoder
-Layout
+Region
 Decoder
 RMSE MRE MAE
-0.3635 0.0571 0.1013
-0.3621 0.0531 0.1011
-0.3656 0.0582 0.1101
-0.3592 0.0511 0.1002
-TABLE VI: The accuracy of resolved background depth maps.
-Methods RMSE MRE MAE
-PanoFormer’s Background Region 0.3214 0.0623 0.1123
-Decode from layout prediction 0.0043 0.0032 0.0023
-Decode from layout ground-truth 0.0001 0.0001 0.0001
-TABLE VII: The comparison of our proposed fusion methods and BGDNet
-Methods RMSE MRE MAE
-Baseline 0.3635 0.0571 0.1013
-Fusion BGDNet 0.2823 0.0498 0.0922
-Fusion ours 0.2436 0.0425 0.0829
-Figure 7: The experiment results of explore the inference of hyper-parameter $\lambda_{2}$ and $\lambda_{3}$.
+0.3017 0.1201 0.1546
+0.3011 0.1109 0.1411
+0.2993 0.1182 0.1424
+0.3012 0.1192 0.1504
+TABLE XII: The training impact of each component to the final depth predictions.
+Background
+Resolving
+Fusion Mask
+Generation
+Adaptive
+Fusion
+RMSE MRE MAE
+0.3017 0.1201 0.1546
+0.2843 0.1011 0.1322
+0.2231 0.0672 0.0833
+0.1935 0.0414 0.0613
+Figure 6: The experiment results of explore the inference of hyper-parameter $\lambda_{2}$, $\lambda_{3}$ and $\lambda_{4}$.
 The x-axis in this figure represents the value of the corresponding hyper-parameter setting, while y-axis represents the RMSE value of our proposed framework’s depth prediction.
-IV-B2 Visualize Comparison with SoTA Depth Estimation Method
-In this section, we perform visualization operations on the scene as shown in the figure.
-In this set of experiments, we selected the baseline panoformer in this paper as the comparison object.
-It is worth noting that in this set of experiments, panoformer and the depth estimation framework proposed in this paper both use denoised datasets during training.
-In this set of experiments, we selected several representative scenes on the validation set for demonstration.
-From the corresponding 3D visualization results, it can be seen that the depth estimated by our method is significantly more accurate in corners and walls.
-At the same time, thanks to the prior knowledge of room layout introduced by us, the 3D visualization of the prediction results of our method is more reasonable in terms of overall structure.
-At the same time, due to the existence of our dataset denoising strategy, panoformer and our method do not have the phenomenon of overfitting of noise points in patches as shown in Fig.
-1.
-The existence of these phenomena fully demonstrates the effectiveness of the depth estimation method we proposed.
-IV-C Visualize and Quatitative analysis of Dataset Denoise Method
-IV-C1 Visualize Analysis of Denoise Method
-We conducted a visualization experiment as shown above for the panoramic depth images of the 2D-3D-S dataset collected in the real world.
-After mapping each pixel in the panoramic depth images to 3D space, we found that there are a large number of noise points in each scene.
-For these noise points, the processing strategy of the existing panoramic depth estimation method is to directly preset a maximum distance threshold.
-In the 2D-3D-S dataset, this threshold is set to 10 meters.
-The depth values of all areas greater than this threshold will be artificially set to 0.
-On the one hand, this processing method destroys the correlation between the pixels of the original panoramic depth map, thereby affecting the learning process of the model.
-On the other hand, it will also cause some areas in the scene with depth values less than 10 but should fall inside the room to fall outside the room.
-In order to preserve the pixel correlation of the processed panoramic depth map and prevent the area falling inside the room from falling outside the room, we proposed a strategy to preprocess the indoor panoramic depth map using pretrained layout.
-In this strategy, we first collected as many parts of the existing dataset that contain layout annotations as possible to form a layout dataset of a relatively reasonable size to pretrain the layout estimation part of the framework proposed in this paper.
-Considering that the layout prediction branch can already obtain a high 3D IoU in our experiments, we directly use this part of the prediction results to process the depth estimation dataset in the actual processing process.
-In our proposed processing strategy, we first use the layout prediction branch to estimate the layout of the current room, and then use the panorama depth ground-truth to count the height of each wall of the room, and use this height result to solve the 2D layout into the corresponding background depth map.
-Finally, we use this background depth map to constrain the original panorama depth ground-truth as shown in Fig.
+In this section, we analyze the impact of each proposed component on the whole.
+Considering that we proposed a framework for auxiliary-task assisted depth estimation tasks, we mainly discuss two aspects of the impact of each component: 1) The impact of hyper-parameter setting for final depth predictions ;2) The impact of auxiliary-task decoder for depth estimation decoder; 3) The impact of each component for the final depth predictions.
+The experimental dataset we chose is Structured3D because it contains complete layout estimation annotations.
+Regarding aspect 1) which we want to discuss, the relevant experimental results are shown in Fig.
 6.
-In this constrained process, we use the background ground-truth as a mask to ensure that the depth of the background region does not exceed the range of the room itself.
-IV-C2 Quatitative Analysis of Denoise Method
-We use the Stanfor2d3d dataset as the object to verify the effectiveness of the proposed noisy image processing strategy.
-In the experiment, we use our proposed framework and PanoFormer to perform training and val on the processed datasets before and after processing.
-The experimental results are shown in Table IV.
-From the results in Table IV, the performance of PanoFormer and our proposed framework trained on the processed dataset is significantly higher than that trained on the original dataset, which fully demonstrates the performance improvement brought by our data processing strategy.
-IV-D Ablation Study
-IV-D1 The Effect of Framework’s Each Component
-Figure 8: The visualize of Background segmentation.
-The left column of the figure is input panorama RGB images, the middle column is background segmentation predictions, the right column is corresponding ground-truth.
-Figure 9: The visualize of room layout estimation.
-The first column shows the input RGB image.
-The second column shows the room layout estimation results displayed on a 2D equirectangular image.
-The top row in each 2D equirectangular sub-image represents the probability of a corner point in that column.
-The third and fourth columns show the corresponding 3D bird’s-eye views of the prediction and layout ground-truth, respectively.
-In this section, we analyze the impact of each proposed component on the whole.
-Considering that we proposed a framework for auxiliary-task assisted depth estimation tasks, we mainly discuss two aspects of the impact of each component: 1) The impact of auxiliary-task decoder for depth estimation decoder; 2) The effect of our proposed background depth resolving and fusion strategy.
-The experimental dataset we chose is Matterport3D because it contains complete layout estimation annotations.
-Regarding the aspect 1), the corresponding experimental results have been put into Table V.
-It is worth noting that the loss weight we used in this group of experiments is the combination verified in subsequent section.
-From the results, it can be seen that the auxiliary-task decoder has very little impact on the depth decoder during training.
-This is primarily because we intentionally set a low loss weight for the auxiliary-task decoder.
-This design stems from two considerations: first, optimizing each task in a multi-task learning framework requires extensive experimentation and manual tuning.
-Second, and more importantly, our framework’s primary objective is to use the auxiliary task’s output to constrain the depth decoder directly, rather than to improve it indirectly by learning more generalized features in the shared encoder.
-Regarding the aspect 2) we want to discuss, we conducted corresponding experiments and put them in Table VI and Table VII.
-The experimental results of the background depth resolving strategy are shown in Table VI, and the experimental results of the fusion strategy used are shown in Table VII.
-When evaluating the background depth resolving strategy, we used the depth ground-truth provided by the Matterport3D dataset as a benchmark.
-When evaluating the resolved background depth, we used the ground-truth of the background segmentation as a mask to ensure that only the background region is evaluated during the eval process.
-Judging from the results in Table VI, our background depth resolving strategy is more accurate than the deep learning model.
-In terms of fusion strategy, we chose BGDNet[32] as the comparison object.
-In this set of experiments, the configuration of the fusion strategy used by BGDNet is consistent with the original paper.
-The experimental results demonstrate that incorporating background depth as a constraint effectively enhances depth estimation performance.
-Moreover, our proposed fusion strategy yields even more substantial improvements compared to that of BGDNet.
+In this set experiments, we divided the weight values of each component in the objective function within the range 0-1 into 0-1 into intervals of 0.1.
+We adopted a component-by-component exploration mode for configuring the optimal loss weights.
+Each time we explored the optimal configuration for a parameter, we set all subsequent parameters to 0 while keeping the preceding parameters unchanged.
+Considering that the layout decoder and other task decoders are trained independently in our framework, the value of its corresponding parameter $\lambda_{1}$ is constant at 0 when training other decoders.
+As shown in Fig.
+6, the framework achieves optimal overall performance when $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$ are set to 1.0, 0.4, and 0.6, respectively.
+Regarding aspect 2), the relevant experimental results are shown in Table XI.
+In this set of experiments, we simply discuss the impact of each task decoder involved in the training process on the depth decoder results.
+In this set of experiments, the Loss weights for each task decoder used the optimal configuration explored in the previous set of experiments.
+As the results shown in the table, different task decoders do not have a significant impact on the final result during training in our framework.
+This also suggests that the parameters explored in the previous set of experiments actually have a greater impact on the final experimental results than on the performance of each task decoder, which in turn affects the performance of the three components we proposed, ultimately impacting the overall framework performance.
+Regarding aspect 3), the relevant experimental results are shown in Table XII.
+In this set of experiments, we used the optimal parameters explored for each task decoder.
+The results in the first row of Table XII.
+represent the experimental results of our baseline Panoformer; the second row represents the performance when we used the proposed background depth resolving component in conjunction with BGDNet’s background replacement strategy; and the third and fourth rows represent the performance when we combined our proposed fusion mask generation and adaptive fusion components, respectively.
+From the results in Table XII., it is clear that the core of our framework’s performance improvement lies in the three task components rather than multi-task learning, and the fusion mask generation component shows the most significant improvement.
+This experimental result also fully validates the core argument of this paper: to make the background depth constraint more reasonable, we need to more carefully identify which regions in the input indoor panorama image belong to the background.
 In summary, our discussion validates both the proposed framework and the collaborative design of its subtasks.
-IV-D2 Analysis of loss weights
-The impact of different loss weights hyper-parameters of the proposed framework has been shown in Fig.
-7.
-Here, we use the MatterPort3D to conduct experiments.
-As the layout decoder in our framework has been pretrained, we simply fix its $\lambda_{1}$ to 0.01.
-For $\lambda_{2}$ and $\lambda_{3}$, we first fix the $\lambda_{3}$ to 1.0 and search the best setting of $\lambda_{2}$.
-Then, fix the $\lambda_{2}$ to the searched results and search the setting of $\lambda_{3}$.
-The Fig.
-6.
-show that on the MatterPort3D dataset, when we set $\lambda_{2}=1.0$ and $\lambda_{3}=0.4$, the proposed framework can achieve the best performance.
-IV-E Visualize Analysis of Background Segmentation and Layout Estimation
-In this section, we visualize the prediction results of the front-segmentation and layout estimation task.
-We used the test set data of Structured3D to perform correspond experiment.
-The specific visualization results are shown in Fig.
-8 and Fig.
-9.
-Fig.
-8 shows the visualization results of the background segmentation task.
-In our framework, the prediction results of the background segmentation task are used as weight to fuse the coarse prediction and background depth.
-When visualizing the results of background segmentation, we mapping the prediction score maps from $[0,1]$ to $[0,255]$ and save them as gray images.
-Judging from the score maps shown in Fig.
-8, our background branch can accurately distinguish between the background and background areas in the input image.
-However, in some areas that are very close to the wall and have a similar depth to the background area, the discrimination ability of our background branch still needs to be improved.
-Fig.
-9 shows the estimation results of our layout estimation branch.
-From the 2D visualization of the prediction results, we can see that the edge areas and corner points of the wall can be identified relatively accurately.
-However, in some areas far away from the camera, the prediction accuracy of the corner points still needs to be improved.
-In addition to these issues, another point worth noting is that the current room layout estimator assumes that the walls of the room are all cubic structures.
-This makes our method perform worse than the normal depth estimator when estimating the depth of some rooms with curved wall structures.
-We consider conducting more detailed research on this issue in future work.
-V Conclusions
-This paper has proposed a panoramic depth estimation framework with room geometry constraints.
-The framework employs multi-task learning, where a shared encoder extracts features that are decoded into three outputs: coarse depth maps, background segmentation masks, and room layouts.
-These predictions enable our method to initially reconstruct background depth from the layout information.
-Subsequently, background depth and coarse depth are fused using the background segmentation mask as a weighting mechanism, ultimately generating the final depth prediction.
-Extensive experiments on real-world and synthetic datasets demonstrate significant performance improvements over current methods.
+V Limitation
+The proposed PAGCNet achieves significantly superior performance on current indoor depth estimation datasets. However, our framework still has several issues:
+1) To adapt to more complex indoor scenes, we divide the scene into regular and irregular regions and only model the background for the regular regions.
+Therefore, we don’t actually model the irregular regions, and naturally, we don’t apply geometric constraints to their depth.
+This makes our framework perform poorly on some extremely irregular room structures, although these scenes are very rare.
+2) In existing panoramic depth estimation datasets, the number of annotations for the room layout task cannot be matched one-to-one with the annotations for tasks like semantic segmentation and depth estimation.
+This mismatch forces us to train the room layout decoder separately before training other decoders.
+Therefore, our framework doesn’t actually solve this problem of annotation imbalance between tasks.
+VI Conclusions
+This paper has proposed a pose-aware and geometry-constrained framework for irregular panoramic depth estimation.Our framework first employs multiple task-specific decoders to jointly estimate room layout, camera pose, depth, and region segmentation from a single input panorama.
+A pose-aware background depth resolving component uses tasks decoder’s prediction to refine the camera pose and subsequently uses the camera pose to compute the background depth of regular enclosed regions, which serves as a strong geometric prior.
+Based on the output of the region segmentation decoder, a fusion mask generation component produces a fusion weight map to guide where and to what extent the geometry-constrained background depth should correct the depth decoder’s prediction.
+Finally, an adaptive fusion component integrates this refined background depth with the initial depth prediction, guided by the fusion weight.
+Extensive experiments on Matterport3D, Structured3D, and Replica datasets demonstrate that our method achieves significantly superior performance compared to current open-source methods.